### v1

# Does Mapo Tofu Contain Coffee? Probing LLMs for Food-related Cultural Knowledge

Abstract
Recent studies have highlighted the presence of cultural biases in Large Language Models (LLMs), yet often lack a robust methodology to dissect these phenomena comprehensively. Our work aims to bridge this gap by delving into the Food domain—a universally relevant yet culturally diverse aspect of human life. We introduce FmLAMA, a multilingual dataset centered on food-related cultural facts and variations in food practices. We analyze LLMs across various architectures and configurations, evaluating their performance in both monolingual and multilingual settings. By leveraging templates in six different languages, we investigate how LLMs interact with language-specific and cultural knowledge. Our findings reveal that (1) LLMs demonstrate a pronounced bias towards food knowledge prevalent in the United States; (2) Incorporating relevant cultural context significantly improves LLMs’ ability to access cultural knowledge; (3) The efficacy of LLMs in capturing cultural nuances is highly dependent on the interplay between the probing language, the specific model architecture, and the cultural context in question. This research underscores the complexity of integrating cultural understanding into LLMs and emphasizes the importance of culturally diverse datasets to mitigate biases and enhance model performance across different cultural domains.
Does Mapo Tofu Contain Coffee?
Probing LLMs for Food-related Cultural Knowledge
[…]
1 Introduction
Figure 1: Summary of the various aspects of our work.
[…]
Trained on vast datasets, Large Language Models (LLMs) encode a wide array of knowledge, but also face challenges with various biases, such as those related to gender Savoldi et al. (2021); Kaneko et al. (2022), belief Søgaard (2021); González et al. (2021); Lent and Søgaard (2021), and culture Cao et al. (2023); Deshpande et al. (2022); Yin et al. (2022). In this paper, we concentrate on probing LLMs for cultural knowledge, aiming to provide evidence of cultural biases inherent in LLMs and explore potential strategies for reducing these cultural biases. Specifically, we begin by introducing the FmLAMA construction method (§3), followed by the knowledge probing method (§4), encompassing the probing template, task, and metric. Subsequently, we conduct experiments employing English prompts on the filtered sub-dataset FmLAMA-en to probe both encoder-only and encoder-decoder LLMs (§5.1). We then analyze various aspects of probing prompt settings (§5.2), including cultural background references, code-switching settings, and language choices within prompts. Following this, we proceed with additional probing experiments targeting decoder-only LLMs (§5.3), considering the prevalence of current advanced LLMs. Finally, we delve into case studies (§5.4) and human evaluations(§6). Figure 1 provides an overview of our work’s various aspects, and our contributions are:
  • We present FmLAMA, a pioneering dataset focused on the food domain, which is inherently rich in cultural diversity Cao et al. (2024). This dataset is a multifaceted tool for probing LLMs across cultures and languages.
  • We propose novel metrics designed to assess LLMs’ ability to accurately and sensitively probe for cultural knowledge, incorporating both absolute- and fuzzy-match techniques.
  • We analyze the impact of integrating cultural context and language specificity in prompts, offering insights to optimize LLMs for equitable cross-cultural knowledge retrieval.
Our methodology for automated collection of cultural knowledge corpora extends the analysis potential in other domains, broadening the scope of research on cultural biases in LLMs.  [^2] Data and code are available in https://github.com/lizhou21/FmLAMA-master.
2 Related Work
Cultural knowledge datasets.
Cultural knowledge, encompassing the customs, beliefs, traditions, and practices of a culture, is crucial yet challenging to encapsulate. While some researchers focus on manually curating cultural knowledge datasets, others evaluate LLMs’ performance on culturally related tasks. Yin et al. (2022) and Palta and Rudinger (2023) have developed benchmarks such as geo-diverse prompts and food-custom datasets (FORK) to probe cultural biases in commonsense reasoning systems. However, manual dataset construction is inefficient and hard to scale, prompting a shift towards automated methods. For instance, StereoKG Deshpande et al. (2022) offers a scalable knowledge graph that blends cultural knowledge with stereotypes, and Candle Nguyen et al. (2023) extracts cultural commonsense knowledge from the web, organizing it into clusters. Despite these advances, the variability in data representation—from sentences to triplets using OpenIE—poses challenges for consistency and noise control in knowledge probing. Keleg and Magdy (2023) aims to mitigate this by selecting culturally diverse factual triples from Wikidata, focusing mainly on explicit country information. In contrast, our work proposes an automated, efficient approach to constructing a cultural knowledge dataset in a uniform triplet format, addressing the limitations of existing methods and focusing on implicit cultural knowledge. Table 1 contrasts our dataset with prior cultural knowledge collections.
Datasets Format Topic Construction method Size
GeoMLama  Yin et al. (2022) Manual Template Geo-Diverse Concept Manually curated 3125
FORK Palta and Rudinger (2023) CommonsenseQA Culinary culture Manually curated 184
StereoKG  Deshpande et al. (2022) Triplet knowledge Stereotypes about religion and ethnicity Automatically constructed 4722
Candle  Nguyen et al. (2023) Sentences Several cultural facets (food, drinks, clothing, traditions, rituals, behaviors) Automatically constructed 47360
FmLAMA (ours) Triplet knowledge Food domain Automatically constructed 33601
Table 1: Comparison of cultural knowledge datasets. Size is in number of instances.
Knowledge probing.
Deciphering the knowledge encoded by LLMs poses significant challenges due to their opaque nature , early benchmarks like LAMA Petroni et al. (2019) sought to quantify the factual knowledge in English LLMs, while ParaRel Elazar et al. (2021) highlighted their consistency issues. Subsequent efforts as mLAMA Kassner et al. (2021) and mParaRel Fierro and Søgaard (2022) expanded these benchmarks multilingually, though such methods often focus on single-word entities, limiting their depth of assessment. To address these shortcomings, newer studies Shin et al. (2020); Zhong et al. (2021); Meng et al. (2022) have evolved towards eliciting more comprehensive factual knowledge, including multi-word entities, with Jiang et al. (2020a) developing algorithms for multi-token predictions. LPAQA Jiang et al. (2020b) further refines this by optimizing prompt discovery for more accurate knowledge probing. Our work builds on this foundation, targeting multi-token probing within the food domain, characterized by complex expressions like Trigonella foenum-graecum and almond paste. We also introduce absolute- and fuzzy-match metrics for a nuanced evaluation of LLMs’ cultural knowledge.
3 FmLAMA Construction
To assess whether LLMs encode and access cultural information, we develop FmLAMA, a multicultural, multilingual dataset focusing on culinary knowledge. The designed framework can be adapted to other cultural domains.
Step #1: Obtain countries set.
Following Zhou et al. (2023), we use countries of food origin to delineate cultural groups. This method leverages countries as proxies for cultural identity, encapsulating diverse traditions, values, and norms that reflect the breadth of human civilizations across geographical boundaries Minkov and Hofstede (2012); Peterson et al. (2018).
Step #2: Acquire food instances.
We utilize SPARQL to query Wikidata, extracting a vast array of food-related data. This approach exploits Wikidata’s RDF triple structure to gather detailed information on food instances, offering a rich source of comprehensive food knowledge.
i. Class. For our food-focused dataset, we concentrate on the dish class and employ two approaches to find food instances:
  • Explicit instance of dish, e.g., bouillabaisse.
[…]
This enables comprehensive inclusion of food instances, represented as $I\xrightarrow{(\text{instance of}|\text{subclass of})+}$ dish, where ‘||||’ denotes “or”, and ‘$+$’ is “one or more”.
ii. Cultural group. We organize food instances by their origin, applying these strategies:
[…]
We exclude dishes with multiple origin countries to maintain cultural specificity.
iii. Properties included. We prioritize the property “has part(s)” to identify food ingredients for each dish. Additional properties like “made from material” and “image” are collected to support future research (e.g., multimodal), though they are not utilized in this study. Language consistency for property descriptions ensures uniformity across the dataset. Our dataset, FmLAMA, comprises 33,601 dish instances, detailed by name, origin, ingredients, and optionally, materials and images. Examples are provided in Appendix A.
Step #3: Filter by language.
To explore various language settings, we further create sub-datasets by applying language filtering to the created FmLAMA dataset. This results in FmLAMA-la, where ‘la’ designates the specific language of the current sub-dataset. In this paper, We focus on a typologically diverse set of languages, namely, English (en), Chinese (zh), Arabic (ar), Korean (ko), Russian (ru), and Hebrew (he), with a filtered sub-dataset of 2590, 815, 571, 807, 961, 462 dishes each. These languages span 4 different language families – Indo-European (English, Russian), Semitic (Hebrew, Arabic), Altaic (Korean), and Sino-Tibetian (Chinese), and are spoken by more than 2.356 billion speakers. Furthermore, these languages represent cultural diversity, being spoken on different continents by groups with rich and distinct cultural backgrounds.
4 Cultural Knowledge Probing
4.1 Probing templates
Following most knowledge-probing methods, we adopt Word Predictions (WP) as knowledge-probing tasks. Specifically, we manually design prompt templates focused on the core attribute “has part(s)”, illustrating a connection between a dish (subject) and its ingredient(s) (object). Considering LLMs produce varied predictions based on prompt framing Elazar et al. (2021); Wang et al. (2023), we craft five templates conveying identical meanings in each prompt language. These templates, in six various involved languages, are depicted in Figure 7 in Appendix B.
Cultural context
To explore the impact of introducing cultural context on LLMs’ ability to access cultural knowledge, we enhance the basic templates by integrating location adverbials. For instance, “In [C], the food [X] includes the ingredients of [Y]”, where [C] denotes the country of origin for the dish [X], [Y] indicates the ingredient object.
Code-switching
To assess the cultural knowledge probing ability of multilingual LLMs under code-switching settings, we configure prompt variations with English-Chinese and Chinese-English CS settings, specifically based on purely English and Chinese prompts, as follows:
  1. EN_en:
    The food beef bourguignon has the ingredients of [Y].
  2. EN_zh:
    The food 红酒炖牛肉 has the ingredients of [Y].
  3. ZH_zh:
    美食红酒炖牛肉是以[Y]为主要原料制作的。
  4. ZH_en:
    美食beef bourguignon是以[Y]为主要原料制作的。
Here, EN and ZH signify the main language of the prompt, while en and zh indicate the subject’s language.
4.2 Probing task
The probing task is defined as a candidate object retrieval problem. The candidate set consists of all objects in the filtered sub-dataset FmLAMA-la. [^3] As the size of the filtered sub-dataset increases, the candidate object set also expands, leading to greater difficulty in probing. Consequently, in this paper, results obtained by probing across different filtered sub-datasets cannot be used for horizontal comparison. Our primary objective is to utilize LLMs to obtain the probability of each candidate $C$ and subsequently rank the predicted objects based on these probabilities.
MASK operation.
We use the corresponding subject-object tuples ([X], [Y]) as the query and probe LLMs by replacing the subject and masking the object. Considering that each candidate object is tokenized into $d$ subtokens $\left\{c_{1},\cdots,c_{d}\right\}$ by LLMs correspondingly, we apply [MASK] token of varying lengths to the objects within each query. So we construct $D$ queries for each food subject $X$ based on the same template $t$, $D$ is the maximum number of object tokens, each query $Q_{d}^{X}$ is defined as:
$Q_{d}^{X}=t\left(X,\left[\mathrm{MASK}\right]*d\right)\vspace{-3pt}.$ (1)
Probability acquisition.
We use Mean Pooling [^4] Mean Pooling is usually used in multi-token probing and much better than max-pooling and first-pooling methods, this is demonstrated by the experimental results of different pooling methods from  Wu et al. (2023). method to obtain the prediction probability of each candidate. Specifically, for candidate object $C=\left\{c_{1},\cdots,c_{d}\right\}$ of length $d$, we obtain its probability from the likelihoods associated with the [MASK] tokens in $Q_{k}^{X}$, and the predicted probability of $C$ is calculated as the average of the probabilities of composing its subtokens:
$P\left(Q_{d}^{X},C\right)=\frac{1}{d}\sum_{i=1}^{d}{p\left(Q_{d}^{X},\left[\mathrm{MASK}\right]_{i}==c_{i}\right)}\vspace{-3pt},$ (2)
where $p\left(\cdot\right)$ is obtained after the log softmax operation.
4.3 Probing Metric
[…]
Mean Average Precision.
mAP is widely used in information retrieval settings, assessing the relevance of predicted objects (in our case ingredients) only when they precisely match the golden ones. The precision at rank $k$ (P@k) for a given food instance $i$ is defined as follows:
$\mathrm{P@k}={\frac{\left|\mathrm{ing}_{i}\cap\mathrm{topk}_{i}\right|}{k}}\vspace{-3pt},$ (3)
where $\mathrm{ing}_{i}$ is the golden ingredients set, $\mathrm{topk}_{i}$ signifies the set of top-$k$ objects with the highest predicted probability of belonging to food item $i$ by LLMs. Then the average precision of food item $i$ is computed as follows:
$\mathrm{AP}_{i}=\frac{1}{\left|\mathrm{ing}_{i}\right|}\sum_{k}^{n}{\mathrm{P}@\mathrm{k}\times\mathrm{rel}@\mathrm{k}}\vspace{-3pt},$ (4)
where $n$ refers to the size of the candidate object set and $\mathrm{rel}@\mathrm{k}$ is a relevance indicator function, which equals 1 if the object at rank $k$ is relevant to food item $i$ and equals to 0 otherwise. Finally, we compute the mAP in the following way:
$\mathrm{mAP}=\frac{1}{\left|G\right|}\sum_{i\in G}{\mathrm{AP}_{i}}\vspace{-3pt},$ (5)
where G represents a food group we are focusing on (i.e. a subset of FmLAMA).
Origin Count Encoder-only LLMs Encoder-Decoder LLMs Avg.
Bb-c Bb-u Bl-c Bl-u mB-c mB-u mT5 T5
Italy 215 (8.3%) 3.93±2.29 5.43±2.26 4.39±2.44 5.32±3.14 5.25±2.40 5.65±4.03 4.49±1.37 5.09±1.82 4.94
U.S. 285 (11.0%) 10.26±4.94 14.58±6.29 11.60±4.67 12.72±6.25 13.79±5.16 13.10±7.62 12.37±3.37 13.16±5.32 12.70
Turkey 98 (3.8%) 7.80±5.13 10.53±5.09 6.72±3.97 8.02±5.66 11.55±6.37 9.12±6.28 6.29±2.77 7.84±3.98 8.48
Japan 186 (7.2%) 7.99±3.75 7.53±4.09 8.68±4.03 6.64±3.75 7.18±3.41 7.59±4.59 5.57±2.63 5.64±1.40 7.10
France 175 (6.8%) 5.01±2.99 6.57±2.90 6.15±2.58 5.55±2.29 5.64±2.62 6.02±3.59 3.63±1.90 4.05±1.67 5.33
U.K. 83 (3.2%) 8.40±5.34 8.84±3.39 11.17±6.01 9.29±5.83 10.48±4.16 8.52±6.48 6.29±3.94 7.71±3.55 8.83
Mexico 57 (2.2%) 6.41±3.12 6.93±2.95 7.72±2.58 7.16±4.55 8.08±2.64 7.92±2.26 3.61±0.97 4.18±2.72 6.50
India 132 (5.1%) 12.63±6.56 14.18±5.99 13.37±7.32 12.10±6.99 9.78±4.34 10.56±5.67 8.64±3.41 5.41±2.55 10.83
Germany 57 (2.2%) 4.94±3.21 5.42±3.39 5.64±3.43 5.78±2.69 7.40±2.65 7.81±4.92 4.20±1.61 5.79±1.26 5.87
China 97 (3.8%) 10.43±4.17 12.35±3.13 12.36±3.34 11.27±4.69 12.54±3.47 11.79±5.54 7.93±1.97 9.36±3.18 11.00
Iran 21 (0.8%) 7.00±5.17 5.96±4.26 6.32±4.56 6.33±3.82 8.31±4.73 9.42±5.85 10.37±3.09 5.67±0.42 7.42
Greece 21 (0.8%) 3.15±2.93 3.99±2.46 2.78±1.84 2.73±1.63 3.72±3.24 3.12±1.42 2.12±0.92 0.73±0.27 2.79
Spain 95 (3.7%) 3.54±1.93 5.92±2.74 4.26±2.24 5.37±2.89 4.61±1.84 5.42±1.90 3.26±1.77 3.01±1.17 4.42
Russia 27 (1.0%) 5.55±3.53 7.68±2.17 5.34±3.21 5.48±3.64 5.31±3.22 5.57±2.58 1.45±1.28 1.85±0.23 4.78
Others 1031 (40.0%) 5.92±3.04 6.30±2.81 6.41±2.88 6.37±3.17 6.06±2.51 5.62±2.42 3.82±1.21 3.61±1.06 5.51
ALL 2580 (100.0%) 6.87±3.46 8.09±3.40 7.59±3.36 7.51±3.77 7.70±2.91 7.43±3.73 5.51±1.75 5.58±1.95 7.04
(a) Performance results evaluated using mAP (%).
Origin Encoder-only LLMs Encoder-Decoder LLMs Avg.
Bb-c Bb-u Bl-c Bl-u mB-c mB-u mT5 T5
Italy 0.3107±0.07 0.3391±0.05 0.3206±0.07 0.3246±0.05 0.2992±0.05 0.3279±0.04 0.3287±0.02 0.3097±0.01 0.3201
U.S. 0.3593±0.08 0.3957±0.07 0.3793±0.07 0.3813±0.07 0.3730±0.05 0.3760±0.06 0.3796±0.03 0.3844±0.04 0.3786
Turkey 0.3431±0.07 0.3592±0.08 0.3542±0.07 0.3521±0.06 0.3478±0.08 0.3480±0.05 0.3501±0.04 0.3243±0.03 0.3474
Japan 0.3247±0.07 0.3321±0.06 0.3312±0.08 0.3072±0.05 0.2902±0.05 0.2965±0.05 0.3072±0.03 0.2815±0.01 0.3088
France 0.3280±0.07 0.3420±0.04 0.3348±0.05 0.3372±0.03 0.3261±0.05 0.3278±0.03 0.3342±0.02 0.3268±0.02 0.3321
U.K. 0.3132±0.08 0.3343±0.03 0.3401±0.06 0.3362±0.05 0.3177±0.04 0.3255±0.05 0.3126±0.04 0.3079±0.02 0.3234
Mexico 0.3123±0.05 0.3418±0.03 0.3473±0.06 0.3254±0.05 0.3087±0.05 0.3152±0.02 0.3276±0.01 0.3130±0.01 0.3239
India 0.3638±0.09 0.3741±0.08 0.3796±0.09 0.3531±0.08 0.3215±0.06 0.3367±0.07 0.3352±0.05 0.2920±0.02 0.3445
Germany 0.3127±0.10 0.3268±0.08 0.3362±0.08 0.3191±0.06 0.3118±0.05 0.3473±0.06 0.3189±0.05 0.3263±0.03 0.3249
China 0.3210±0.04 0.3493±0.02 0.3476±0.02 0.3354±0.03 0.3398±0.02 0.3309±0.04 0.3569±0.02 0.3156±0.03 0.3371
Iran 0.3195±0.07 0.3145±0.09 0.3399±0.08 0.3145±0.08 0.3252±0.07 0.3476±0.04 0.3555±0.03 0.3233±0.04 0.3300
Greece 0.3113±0.08 0.3399±0.06 0.3359±0.08 0.3178±0.06 0.2822±0.08 0.3180±0.02 0.3530±0.05 0.3048±0.01 0.3204
Spain 0.3008±0.06 0.3280±0.05 0.3189±0.06 0.3117±0.04 0.2856±0.05 0.3068±0.03 0.3168±0.02 0.2828±0.01 0.3064
Russia 0.3368±0.05 0.3680±0.04 0.3355±0.06 0.3397±0.04 0.3113±0.07 0.3158±0.02 0.3313±0.03 0.2721±0.02 0.3263
Others 0.3235±0.06 0.3378±0.04 0.3376±0.06 0.3263±0.04 0.3088±0.05 0.3150±0.02 0.3266±0.02 0.2821±0.02 0.3197
ALL 0.3278±0.06 0.3468±0.05 0.3428±0.06 0.3339±0.05 0.3177±0.05 0.3264±0.03 0.3338±0.02 0.3050±0.02 0.3293
(b) Performance results evaluated using mWS.
Table 2: Probing performance comparison with English prompts and FmLAMA-en sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT. “b/l” stands for base/large and “c/u” stands for cased/uncased. Bold represents the best-performing cultural group within the same model (each column). Red indicates the best-performing LLMs in each specific cultural group (each row). We find that LLMs typically score higher on U.S. cultural knowledge, and monolingual English LLMs perform better in English-speaking countries. The Pearson correlation coefficient between mAP and mWS is 0.66 based on the results above.
Mean Word Similarity.
mWS is defined based on the semantic similarity between predicted and golden objects. First, we define the similarity score $S\left(i,g\right)$ for each ingredient $g$ within each food instance $i$. Only the predicted objects in the top-$l$ rankings that are most similar to $g$ contribute to the evaluation score, where $l$ is the size of $\mathrm{ing}_{i}$. Mathematically, $S\left(i,g\right)$ is defined as follows:
$S@l\left(i,g\right)=\underset{p\in\mathrm{topl}_{i}}{\max}\left[\cos\left(w_{g},w_{p}\right)\right],g\in\mathrm{ing}_{i}\vspace{-3pt},$ (6)
where $w_{g}$ and $w_{p}$ are the continuous word representations for the ingredient $g$ and the predicted object $p$, respectively, and are obtained using Fasttext Bojanowski et al. (2017); Joulin et al. (2016), $\cos\left(\cdot\right)$ is the cosine similarity function. Then we can compute the probing similarity $WS_{i}$ for each food instance $i$ and mWS for the targeted food group as follows:
$\mathrm{WS}_{i}=\frac{1}{\left|\mathrm{ing}_{i}\right|}\sum_{g\in\mathrm{ing}_{i}}{S@l\left(i,g\right)}$ (7)
$\mathrm{mWS}=\frac{1}{\left|G\right|}\sum_{i\in G}{\mathrm{WS}_{i}}$ (8)
Both mAP and mWS metrics range from [0,1].
5 Experiments
5.1 Access evaluation
We explore the masked language models, including the encoder-only language model BERT Devlin et al. (2019) and its multilingual version, mBERT, along with the encoder-decoder language model T5 Raffel et al. (2020) and its multilingual counterpart, mT5 Xue et al. (2020). Table 2 showcase the probing results based on English prompt on the filtered dataset FmLAMA-en. [^5] The probing results with prompts in the other five languages on the corresponding filtered sub-datasets can be found in the Appendix C.
Metric comparison
Regarding the adopted two metrics, we observe the following: 1) Across diverse cultural groups, evaluating probing results with the absolute-match metric, mAP, reveals a significantly greater disparity compared to assessments using the fuzzy-match metric, mWS. This highlights the challenge of absolute-match evaluation and the importance of introducing the fuzzy-match metric. 2) We evaluate the correlation between the two metrics, mAP, and mWS, utilizing all probing results from Table 2, resulting in a Pearson correlation coefficient of 0.66. This signifies a moderate to strong positive correlation between the two metrics.
Group comparison
Regarding the various cultural groups in Table 2, we observe the following: (1) Regardless of absolute or fuzzy evaluation, the U.S. group almost consistently achieves the highest probing results on FmLAMA-en with English prompts. This suggests that evaluated LLMs possess a greater familiarity with food-related knowledge specific to the U.S. context. (2) The quantity of knowledge within cultural groups in the knowledge base may not fully reflect the potential knowledge within LLMs. For example, despite the Italy group having more dishes than the U.K. and India groups, LLMs achieve lower probing result scores for the Italy group. We speculate that this difference may be due to the official languages of the U.K. and India including English, aligning with the prompt language in the experiments. However, it is notable that LLMs do not exhibit lower probing result scores in China, despite its official language not including English. (3) Monolingual English LLMs excel in accessing cultural knowledge within English-speaking countries, while multilingual LLMs may not demonstrate superior performance in non-English-speaking countries. This discrepancy may arise from the fact that the pretraining data for monolingual models is solely in English, whereas for multilingual models, it encompasses languages beyond English.
Figure 2: Comparative probing results on FmLAMA-en: incorporating cultural information about the origin of dishes into English prompts can enhance the probing of cultural knowledge.
5.2 Prompt analysis
In this part, we examine the impact of different prompt settings on cross-cultural knowledge exploration. This involves integrating references to cultural backgrounds, incorporating code-switching settings, and considering language choices within the prompts.
Cultural background analysis
Using the English prompts from Figure 7 as a basis, we incorporate information about the country of origin into the probing prompts for each specific dish (as described in §4.1). Figure 2 presents a comparative analysis of probing results on FmLAMA-en across different LLMs, taking into account the inclusion of cultural information mentions in the prompts. We find that, whether through absolute or fuzzy evaluation, English prompts with cultural information achieve higher probing scores in capturing LLMs’ knowledge within the food domain. This suggests the importance of emphasizing the cultural background when utilizing LLMs, especially in the exploration of culture-related topics. Moreover, in the comparison of the absolute-match metric mAP, cultural information contributes more in multilingual LLMs compared to monolingual LLMs. This seems to align with real-world social dynamics, indicating that in multicultural settings, introducing cultural backgrounds enhances communication.
Figure 3: Probing results with monolingual prompts and code-switching prompts, $\circ$ and $\Box$ signify the absence or presence of cultural background information in the prompt, respectively.
Code-switching analysis
Code-switching probing focused on dishes common in both English and Chinese in a filtered subset of FmLAMA. Only cultural groups with dish counts exceeding 20 are retained. Our objective is to ensure that the language of the predicted ingredients matches the main language used in the prompt. Probing results on each multilingual LLMs display for Figure 3: (1) In code-switching settings, the probing abilities of LLMs for EN_zh relative to EN_en and ZH_en relative to ZH_zh both notably decrease, with the latter showing a more pronounced decline in the mB-c language model. (2) Whether in code-switching mode or not, the introduction of relevant cultural backgrounds in the prompts generally facilitates cultural knowledge probing. The detailed probing results of each cultural group can be seen in Appendix D.
Figure 4: Average probing results of all multilingual LLMs on the filtered sub-dataset with prompts in different languages. The ability of LLMs to access cultural knowledge may not consistently outperform in the cultural group where that language is spoken for each language prompt.
Language analysis
We adopt prompts in six different languages and conduct knowledge probing on multilingual LLMs. To ensure a fair comparison of the probing results across different language prompts, we filter a sub-dataset consisting of only 175 food instances that have labels in all of the languages that are involved. This underscores the lack of aligned knowledge across multiple languages. Considering the relatively limited size of the filtered dataset, we opt for a broader categorization based on continents rather than individual countries to differentiate cultural groups. Appendix E illustrates the distribution of cultural groups within this filtered sub-dataset. Figure 4 displays probing results on the filtered sub-dataset with prompts in different languages. The probing performance varies with prompts in different languages. The ability of LLMs to access cultural knowledge may not consistently outperform in the cultural group where that language is spoken for each language prompt. For instance, Chinese prompts yield better results in English-speaking countries. Our experimental results suggest that determining the most suitable language for knowledge probing in a specific cultural context proves challenging.
Prompt language EN ZH
Subject language en zh en zh
Llama2-7b-chat-hf 27.36 17.91 7.28 5.91
Llama2-13b-chat-hf 35.04 22.64 10.43 7.87
vicuna-7b 36.22 19.69 12.01 14.37
vicuna-13b 28.15 20.87 17.72 16.54
gpt-3.5-turbo 47.64 34.06 20.28 25.00
Table 3: Accuracy (%) of decoder-only LLMs’ results.
5.3 Decoder-only LLMs probing
LLMs, particularly decoder-only models, offer robust capabilities in both natural language understanding and generation. We specifically conduct cultural knowledge probing on instruction-tuned models, including Llama2-chat Touvron et al. (2023), Vicuna Chiang et al. (2023), and gpt-3.5-turbo OpenAI (2023). Considering the characteristics of these LLMs, we employ a naive approach to probe and evaluate them. [^6] The experimental setup details, evaluation method, cultural knowledge probing results for each cultural group, and additional analysis are provided in Appendix F. Table 3 shows the overall probing results using accuracy for both the monolingual and code-switching prompt settings. We observe that gpt-3.5-turbo demonstrates the highest extent of cultural knowledge. Additionally, all decoder-only LLMs demonstrate better performance with monolingual English prompts compared to monolingual Chinese prompts, and the performance with code-switching prompts is inferior to that with their corresponding monolingual prompts.
Most common ingredients Gold label ingredients
en zh ko en zh ko
yogurt water tangerine egg grain flour egg
avocado sugar cumin seeds flour egg wheat flour
vegetable glucose hot stone sugar sugar sugar
olive oil oil lime juice potato rice potato
rice salt pork floss almond chicken salt
bread egg rice meat tomato chicken meat
baking powder quince drinking water onion onion butter
ice cream milk apple tomato butter meat
extra virgin olive oil sugar cheese chicken meat pork milk
vegetable oil cilantro tofu butter potato bread
Table 4: Ingredients analysis. The first three columns display the 10 most common ingredients for mBERT-base-uncased in English (en), Chinese (zh), and Korean (ko). The remaining three columns show the 10 most common gold standard ingredients, sourced from Wikipedia.
5.4 Case study – Ingredients analysis
We conduct a more fine-grained analysis of the model’s predictions to better understand its behavior, particularly to discern whether it is relying on “educated” guesses or leveraging its cultural knowledge when providing answers. We focus on two multilingual models, mT5 and mbert-base-uncased, and three languages: Korean, Chinese, and English [^7] This subset of languages is selected for clarity. The observed trends are consistent across all languages in our dataset.. For every language and dish in the multilingual dataset (see §3), we extract the top 5 ingredients with the highest probability as predicted by the model and calculate the set of distinct ingredients. Table 4 shows mbert’s 10 most common ingredients in each language. In Korean, Chinese, and English, these ingredients cover $0.59\%$, $0.91\%$, and $0.97\%$ of the total predictions, respectively. [^8] To compute the coverage we divide the number of times the most common 10 predictions are predicted by the model, by the number of its top 5 predictions. For each dish we take only 5 predictions because more than $0.95\%$ of the dishes contain less than 5 ingredients. This indicates that the model tends to depend on repetitive predictions, consistently selecting the same subset of ingredients regardless of the context.  [^9] This trend persist for all models, and languages. For example, in Korean, mbert often predicts ingredients specific to Korean dishes, such as hot stone, which refers to various Korean dishes served in a hot stone pot, and tofu. However, in Chinese/English, the model tends to predict more widely used ingredients, like flour and sugar, which are common across various cultural cuisines. This observation might imply that Chinese/English models perform better because the ingredients they predict have broader applicability and are more likely to be found in a variety of dishes.
6 Human evaluation
To check the validity of the two proposed metrics, we conduct a human evaluation of the probing results, including 372 food instances. Specifically, we extract the probing results for all dishes originated in the U.S and China from the best-performing prompt setting in the code-switching EN_en experiments. In each dish instance, only the top-$l$ predicted objects are evaluated, where $l$ is the number of golden labels, similar to the setting of the mWS metric. In addition to assessing the absolute matching between predicted objects and golden objects as measured by mAP, our human evaluation also considers: (1) whether a predicted object can replace a certain golden object in cooking, and (2) whether the predicted object is an ingredient of the dish, even if not listed in the golden label (e.g, lemon is an ingredient of the dish lemon chicken, even though it erroneously has only chicken meat in Wikidata).
The calculation method for human evaluation scores is the same as mAP. Four authors participated in human evaluation for this work, who come from three different countries, three of whom have experience living across various cultures. The Krippendorff’s alpha value about inter-annotator agreement is 0.581, which suggests a moderate level of agreement among raters. We average their final scores to obtain the ultimate assessment score. Table 5 compares the scores from human evaluation with those from two automated metric evaluations. The Pearson correlation coefficient between mAP and Human is 0.88, whereas that between mWS and Human is 0.94 as per the aforementioned results, which indicates the significance of mWS, as it aligns more closely with human evaluation.
Origin mBERT-uncased mT5
mAP mWS Human mAP mWS Human
U.S. 10.68 37.96 33.61 14.70 38.24 37.56
China 8.40 32.79 25.16 9.40 37.41 32.19
Table 5: Comparison of scores (%) from human evaluation and two automated metric evaluations. The Pearson correlation coefficient between mAP and human is 0.88, whereas that between mWS and human is 0.94 as per the aforementioned results.
7 Conclusion
This study presents an automated method for generating extensive cultural knowledge datasets, exemplified by the creation of FmLAMA, a diverse, food-centric dataset that spans multiple cultures and languages. We introduce novel metrics for cultural knowledge evaluation in LLMs, emphasizing the influence of cultural context and language in the probing process. Our findings reveal a predominant bias towards American culture in LLMs when using English prompts, a bias that diminishes with prompts in other languages. Interestingly, incorporating explicit cultural cues in prompts enhances LLMs’ cultural knowledge access. The study also highlights the scarcity of culturally diverse knowledge across languages, pointing to a potential root of observed biases in LLMs.
8 Limitations
While this study provides valuable insights into cross-cultural knowledge probing in LLMs, it is essential to acknowledge several limitations. Firstly, the food domain knowledge dataset utilized in this research is sourced from Wikidata, which may not offer comprehensive coverage. For example, the dish soy sauce chicken may only include the ingredient chicken meat while lacking the inclusion of soy sauce. Moreover, ingredient descriptions are not always detailed. For instance, the Wikidata gold label might be oil when the recipe requires a specific type of oil, such as sesame oil. This inconsistency underscores the motivation behind our mWS metric. Furthermore, aside from well-known dishes, certain recipes lack standardization and may vary depending on individual preferences and cooking styles, posing challenges to precise probing. Additionally, the fuzzy-match metric mWS, introduced in this study, relies on Fasttext for obtaining object representation vectors. However, for certain objects in Chinese and Korean, zero vectors may result, rendering similarity calculation impossible. Lastly, we employ manually crafted templates in this paper. However, research has shown that sampling templates from large corpora can also enhance knowledge-probing evaluation. This aspect is deferred to future work. Despite our endeavors to construct comprehensive multilingual and multicultural knowledge repositories, the availability of aligned cross-cultural knowledge remains limited in multilingual settings. This constraint presents challenges in exploring the interaction between language and culture.
Appendix A FmLAMA examples
Figure 6 illustrates examples from dataset FmLAMA. Each dish instance in FmLAMA is defined as: $\left(url,na,cou,la,pa,\left[ma,im\right]\right)$, the elements in []delimited-[][$\cdot$][ ⋅ ] indicate optional:
  • $url$: the link in Wikidata;
  • $na$: the name of the dish;
  • $cou$: the country of origin of the dish;
  • $la$: the language used in this entry;
  • $pa$: the ingredients of this dish;
  • $ma$: the material used in the dish;
  • $im$: the image of this dish.
Particularly, in each instance, $la=\mathrm{LANG}\left(na\right)=\mathrm{LANG}\left(pa\right)=\mathrm{LANG}\left(ma\right)$. For a dish with the same $url$, there may be several instances with different languages. There are totally of 33,601 entries in FmLAMA.
Appendix B Probing Templates
Probing templates in six involved languages are shown in Figure 7, in which [X] represents the subject (dish) and [Y] indicates the object (ingredient).
Appendix C Probing results in each FmLAMA-la
Figure 5: Probing results with code-switching settings in each cultural group, w/o and w/ signify the absence or presence of cultural background information in the prompt, respectively.
Figure 6: Examples of FmLAMA.
Figure 7: Probing templates in six involved languages, with [X] representing the subject and [Y] indicating the object that can be substituted.
Besides the multilingual LLMs discussed in §5.1, we also probe XLM-RoBERTa Conneau et al. (2020) here. [^10] Indeed, we probe XLM-RoBERTa on FmLAMA-en as well, yielding very poor probing results. Nonetheless, although XLM-RoBERTa’s performance remains subpar in other languages, it shows some relatively positive outcomes. Furthermore, in addition to Hebrew, we configure probing for monolingual LLMs in Arabic, Russian, Korean, and Chinese, including asafaya/bert-base-arabic Safaya et al. (2020), DeepPavlov/rubert-base-cased Kuratov and Arkhipov (2019), kykim/bert-kor-base, klue/bert-base Park et al. (2021) and bert-base-chinese.
The probing results with prompts in the other five languages (Arabic, Hebrew, Russian, Korean, and Chinese) on the corresponding filtered sub-datasets (FmLAMA-ar, FmLAMA-he, FmLAMA-ru, FmLAMA-ko, and FmLAMA-zh) are depicted in Table 6 through Table 10, respectively. Because certain objects in Chinese and Korean have representation vectors that result in all zeros when obtained through Fasttext, calculating cosine similarity was not feasible. Consequently, mWS evaluation was not conducted for prompts in Chinese and Korean prompts. Overall, irrespective of the language used for probing, LLMs still exhibit a relatively strong familiarity with knowledge in the food domain within American culture. However, they may show a slight preference for certain cultural knowledge; for example, when probing in Arabic, LLMs may show a better capability in probing knowledge related to Iranian groups. Additionally, the probing capability of monolingual LLMs may not necessarily surpass that of multilingual LLMs.
Appendix D Cultural results for Code-switching probing
For a more detailed comparison of the knowledge probing abilities of LLMs across various cultural groups, we evaluate the probing results for each cultural group individually. Figure 5 shows the average probing results for the three multilingual LLMs in each cultural group, from which we find: (1) English-dominant prompt settings generally outperform those where Chinese is the primary language. (2) En_en prompt setting excels in probing food knowledge for U.K., U.S., India, and China cultures, while proficiency in the ZH_zh prompt setting is also notable for the above cultural groups except India. (3) In code-switching settings, the detection abilities of EN_zh relative to EN_en and ZH_en relative to ZH_zh both notably decrease, with the latter showing a more pronounced decline. (4) Whether in code-switching mode or not, introducing relevant cultural backgrounds in the prompts aids in the detection of cross-cultural knowledge, manifested in overall probing results and the probing ability within almost every cultural group. The more detailed probing results for each multilingual LLM across various cultural groups can be found in Tables 11 and 12.
Appendix E Data distribution of 175 dishes
The figure 8 illustrates the data distribution of the filtered sub-dataset used in the language analysis in §5.2. It encompasses both continent-level and country-level data distributions.
(a) Continent-level distribution
(b) Country-level distribution
Figure 8: Data distribution for the filtered sub-dataset comprising 175 dishes.
Appendix F Decoder-only LLM probing details
In this section, We present the experimental setup for knowledge probing with decoder-only LLMs, and showcase their fine-grained probing results across various cultural groups.
F.1 Experiential setup
Decoder-only LLMs
: LlaMa2-chat Touvron et al. (2023), a collection of pre-trained and fine-tuned chat models; Vicuna Chiang et al. (2023), a chat assistant trained by fine-tuning LLaMA on user-shared conversations collected from ShareGPT; and gpt-3.5-turbo* OpenAI (2023), a GPT-3.5 model fine-tuned on human instructions using Reinforcement Learning with Human Feedback (RLHF).
Prompt construction
Decoder-only LLMs, in contrast to encoder-only LLMs and encoder-decoder LLMs, focus solely on generating text based on contextual information provided to them. Considering these factors, the probing task involving predicting the probability of the [MASK] token during decoder-only LLM probing will no longer be applicable. We extend the initial template with fill-in instructions, such as “The food [X] has the ingredients of []. Please fill in the sentence.” We only select the best-performing template for extension and probing of the decoder-only LLMs here. Taking into account the diversity of responses from chat LLMs and the ease of evaluating the probing results, we also consider employing an alternative fill-in instruction: “Please complete the sentence with only one entity”. The final results only showcase the probing outcomes under the best-performing settings.
Evaluation
For each dish instance, we parse the text generated by the decoder-only LLMs to obtain the final predicted object list. This object list is not derived from a candidate object set, resulting in greater diversity. Therefore, before conducting the matching evaluation, we lemmatize each object in both the golden object list and the predicted object list to obtain the lemmatized format of each word, such as potatoes $\rightarrow$ potato. We utilize accuracy (ACC) as the metric here, employing a forgiving form of absolute matching to assess result accuracy. Specifically, for each dish instance, if any predicted object matches any object in the golden object list, we consider the prediction for that dish to be correct.
F.2 Cultural probing results
Table 13 show the probing results in each cultural group, including monolingual prompts and code-switching prompts. In monolingual English prompts, all decoder-only LLMs exhibit superior probing performance for knowledge related to the cultural groups of Italy, the U.S., and France. When using English-Chinese code-switching prompts, the overall probing performance tends to decrease, but the best-performing cultural groups remain largely unchanged. However, there are some improvements in the probing rankings for the China group in Llama2-7b-chat-hf and the U.K. group in vicuna-13b. In the case of monolingual Chinese prompts, the cultural groups with the best probing performance are primarily concentrated in China, the U.K., and the U.S. Similarly, when using Chinese-English code-switching prompts, the overall probing performance tends to decrease.
Origin Count Bb-ar Bl-ar mB-c mB-u XRb XRl mT5 Avg.
Italy 76 (13.3%) 2.16±0.95 2.55±0.97 3.52±1.73 2.55±1.25 1.33±0.63 1.30±0.43 3.35±0.71 2.39
U.S. 54 (9.5%) 3.07±1.45 6.01±3.35 4.74±2.46 4.32±4.47 2.88±2.46 2.26±1.05 6.17±0.97 4.21
Turkey 52 (9.1%) 4.59±2.30 5.97±4.08 4.58±1.37 4.41±3.98 2.50±1.14 1.93±0.11 3.18±0.90 3.88
Japan 44 (7.7%) 3.02±1.11 2.86±1.01 3.76±1.54 2.80±0.56 1.69±0.93 1.30±0.10 2.18±0.77 2.52
France 37 (6.5%) 5.25±2.06 5.87±4.09 4.92±3.30 3.98±3.64 4.28±4.26 3.67±2.76 6.17±2.40 4.88
U.K. 26 (4.6%) 6.78±2.88 8.97±6.33 8.53±5.48 6.02±6.67 2.77±3.28 2.18±1.86 8.84±5.02 6.30
Mexico 20 (3.5%) 2.95±0.40 5.44±2.22 4.52±2.85 5.74±1.82 2.08±0.34 1.88±0.27 3.24±1.27 3.69
India 18 (3.2%) 3.67±3.17 6.41±3.58 9.02±7.34 4.44±5.49 1.90±0.27 1.79±0.13 5.55±2.93 4.68
Germany 13 (2.3%) 6.92±3.83 9.67±5.54 8.03±4.78 6.05±4.50 1.45±0.72 1.49±0.65 5.41±3.72 5.57
China 13 (2.3%) 3.25±2.55 2.91±2.51 4.44±3.22 4.33±4.76 1.93±0.16 1.76±0.10 4.04±2.01 3.24
Iran 11 (1.9%) 7.74±5.56 11.09±5.02 6.61±4.79 4.61±5.41 4.04±4.51 4.10±4.54 6.02±4.34 6.32
Greece 11 (1.9%) 3.31±0.85 4.98±4.35 4.16±3.98 4.66±4.18 1.79±1.15 1.72±0.89 4.04±1.80 3.52
Spain 10 (1.8%) 6.03±2.68 4.75±3.35 6.16±4.09 3.04±2.87 3.46±1.32 2.79±0.27 5.53±1.39 4.54
Russia 10 (1.8%) 3.36±0.63 2.32±0.49 1.93±0.27 1.82±0.32 2.05±0.41 1.91±0.21 2.47±0.48 2.27
Others 176 (30.8%) 4.03±0.97 3.54±1.57 4.30±1.40 2.83±1.72 2.74±1.19 2.31±0.15 4.56±1.59 3.47
ALL 571 (100.0%) 3.95±1.31 4.66±2.36 4.73±2.03 3.61±2.66 2.48±1.46 2.10±0.59 4.53±1.37 3.72
(a) Performance results evaluated on mAP (%).
Bb-ar Bl-ar mB-c mB-u XRb XRl mT5 Average
Italy 0.3147±0.02 0.2786±0.05 0.2917±0.04 0.2383±0.02 0.1991±0.02 0.1930±0.00 0.3120±0.03 0.2611
U.S. 0.2943±0.03 0.3057±0.05 0.3087±0.02 0.2648±0.05 0.2208±0.05 0.2072±0.02 0.3264±0.02 0.2754
Turkey 0.2996±0.03 0.2754±0.07 0.2787±0.04 0.2432±0.03 0.1941±0.03 0.1866±0.01 0.2867±0.02 0.2520
Japan 0.2977±0.02 0.2416±0.04 0.2685±0.02 0.2409±0.03 0.1795±0.02 0.1729±0.01 0.2791±0.03 0.2400
France 0.3360±0.02 0.3221±0.05 0.3032±0.04 0.2877±0.03 0.2255±0.08 0.2103±0.05 0.3333±0.04 0.2883
U.K. 0.3125±0.05 0.3418±0.06 0.3632±0.06 0.2736±0.09 0.2270±0.07 0.2103±0.03 0.3681±0.04 0.2995
Mexico 0.3273±0.02 0.2866±0.04 0.3036±0.03 0.2923±0.04 0.2039±0.04 0.2043±0.04 0.3036±0.04 0.2745
India 0.2707±0.03 0.2726±0.07 0.3030±0.04 0.2261±0.06 0.1950±0.02 0.1814±0.01 0.2808±0.03 0.2471
Germany 0.3106±0.03 0.3321±0.06 0.3275±0.05 0.2758±0.05 0.2048±0.03 0.1923±0.00 0.3113±0.02 0.2792
China 0.3006±0.04 0.2658±0.04 0.3366±0.03 0.2854±0.06 0.2035±0.02 0.2000±0.01 0.3037±0.05 0.2708
Iran 0.3714±0.05 0.3420±0.06 0.3370±0.05 0.2687±0.05 0.2291±0.07 0.2286±0.07 0.3289±0.04 0.3008
Greece 0.3068±0.01 0.3067±0.06 0.3260±0.05 0.2693±0.03 0.2270±0.01 0.2306±0.02 0.3147±0.04 0.2830
Spain 0.3930±0.03 0.3049±0.06 0.3487±0.06 0.3124±0.03 0.2377±0.03 0.2395±0.03 0.3418±0.05 0.3111
Russia 0.3353±0.01 0.3068±0.04 0.2885±0.03 0.2784±0.01 0.2264±0.05 0.2183±0.03 0.3157±0.03 0.2813
Others 0.2989±0.01 0.2720±0.05 0.2932±0.04 0.2452±0.01 0.2049±0.04 0.1939±0.01 0.3001±0.02 0.2583
ALL 0.3078±0.02 0.2854±0.05 0.2999±0.03 0.2549±0.02 0.2065±0.04 0.1972±0.02 0.3083±0.02 0.2657
(b) Performance results evaluated on mWS.
Table 6: Probing performance comparison with Arabic prompts and FmLAMA-ar sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT, “XR” denotes XLM-RoBERTa, “b/l” stands for base/large and “c/u” stands for cased/uncased. Bold and underline represent the best-performing and second-performing cultural group within the same model. The Pearson correlation coefficient between mAP and mWS is 0.71 based on the results above.
Origin Count mB-c mB-u XRb XRl mT5 Avg.
Italy 59 (12.7%) 5.73±2.00 3.62±1.43 3.60±0.16 4.04±0.85 4.72±0.60 4.34
U.S. 67 (14.5%) 11.59±2.64 6.72±2.59 4.51±0.81 5.05±0.58 11.55±2.31 7.88
Turkey 12 (2.6%) 9.29±7.33 9.31±5.64 8.83±0.45 9.00±0.78 13.64±7.43 10.01
Japan 19 (4.1%) 8.37±4.61 5.81±3.26 7.39±1.97 8.02±0.77 8.56±3.44 7.63
France 65 (14.0%) 9.26±1.15 3.55±1.25 3.61±0.25 4.03±0.54 4.63±1.26 5.02
U.K. 23 (5.0%) 16.17±3.36 9.89±4.31 3.54±0.78 3.43±0.84 8.02±1.97 8.21
Mexico 11 (2.4%) 2.38±0.57 2.30±0.86 4.90±1.97 4.93±1.48 2.59±1.09 3.42
India 18 (3.9%) 4.83±3.24 5.68±6.31 5.30±2.41 6.45±0.15 6.30±4.71 5.71
Germany 11 (2.4%) 15.41±5.12 8.46±4.91 2.97±0.33 3.67±1.64 14.20±4.70 8.94
China 9 (1.9%) 3.94±3.35 2.63±1.49 6.12±1.36 6.47±0.97 4.19±4.65 4.67
Iran 6 (1.3%) 16.57±7.01 7.54±6.06 6.15±2.83 6.27±2.68 6.49±4.44 8.60
Greece 6 (1.3%) 4.33±2.31 1.85±0.60 6.73±2.89 6.82±2.87 4.24±1.18 4.79
Spain 7 (1.5%) 5.17±2.75 2.74±0.72 5.54±2.15 5.77±1.56 7.62±3.32 5.37
Russia 4 (0.9%) 5.22±2.97 4.82±3.18 5.91±0.27 5.44±0.84 3.53±1.26 4.98
Others 146 (31.5%) 5.48±0.57 3.60±1.46 4.94±0.84 5.15±0.43 4.38±1.34 4.71
ALL 463 (100.0%) 7.90±1.15 4.77±1.90 4.70±0.73 5.05±0.08 6.42±1.33 5.77
(a) Performance results evaluated on mAP (%).
mB-c mB-u XRb XRl mT5 Average
Italy 0.3313±0.01 0.3286±0.03 0.2963±0.00 0.3032±0.02 0.2987±0.07 0.3116
U.S. 0.3706±0.02 0.3442±0.04 0.3090±0.01 0.3130±0.02 0.3565±0.07 0.3387
Turkey 0.3085±0.07 0.3182±0.05 0.3048±0.04 0.2958±0.01 0.2868±0.15 0.3028
Japan 0.2963±0.04 0.2998±0.06 0.2765±0.02 0.2708±0.01 0.2868±0.09 0.2860
France 0.3429±0.01 0.3278±0.04 0.2987±0.02 0.2948±0.01 0.3089±0.05 0.3146
U.K. 0.3832±0.04 0.3776±0.02 0.2777±0.03 0.2681±0.01 0.3217±0.06 0.3257
Mexico 0.3585±0.02 0.3534±0.09 0.3444±0.03 0.3260±0.01 0.3605±0.06 0.3486
India 0.3401±0.03 0.3431±0.06 0.2809±0.02 0.2901±0.04 0.2738±0.13 0.3056
Germany 0.3703±0.06 0.3536±0.04 0.2923±0.06 0.2829±0.04 0.3623±0.07 0.3323
China 0.2713±0.02 0.3245±0.05 0.2646±0.04 0.2531±0.02 0.2541±0.12 0.2735
Iran 0.3856±0.08 0.3494±0.08 0.3827±0.00 0.3642±0.04 0.3724±0.06 0.3709
Greece 0.3310±0.02 0.3040±0.02 0.3393±0.02 0.3388±0.02 0.3286±0.05 0.3283
Spain 0.3128±0.03 0.3768±0.06 0.3646±0.01 0.3488±0.03 0.4115±0.10 0.3629
Russia 0.3431±0.04 0.3713±0.07 0.3803±0.04 0.3596±0.00 0.3268±0.06 0.3562
Others 0.3048±0.01 0.3105±0.05 0.2890±0.02 0.2842±0.01 0.2782±0.08 0.2933
ALL 0.3321±0.01 0.3287±0.04 0.2979±0.02 0.2952±0.01 0.3068±0.07 0.3121
(b) Performance results evaluated on mWS.
Table 7: Probing performance comparison with Hebrew prompts and FmLAMA-he sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT, “XR” denotes XLM-RoBERTa, “b/l” stands for base/large and “c/u” stands for cased/uncased. Bold and underline represent the best-performing and second-performing cultural group within the same model. The Pearson correlation coefficient between mAP and mWS is 0.31 based on the results above.
Origin Count Bb-ru mB-c mB-u XRb XRl mT5 Avg.
Italy 101 (10.5%) 2.97±2.52 6.28±3.53 8.19±3.72 1.79±0.13 1.61±0.29 4.35±1.19 4.20
U.S. 79 (8.2%) 6.86±2.93 8.24±3.97 11.05±3.62 2.29±0.82 1.81±0.15 5.49±1.63 5.96
Turkey 28 (2.9%) 4.12±3.32 4.19±2.62 7.92±0.85 1.98±0.06 1.82±0.50 3.82±3.07 3.98
Japan 68 (7.1%) 4.41±2.32 3.79±1.22 5.37±2.72 1.67±0.22 1.52±0.02 1.91±1.15 3.11
France 93 (9.7%) 4.36±2.40 3.92±2.44 4.82±3.74 1.91±0.67 1.52±0.16 4.86±3.00 3.56
U.K. 33 (3.4%) 6.03±6.40 7.81±3.94 9.74±3.14 3.14±3.90 1.37±0.19 9.79±6.82 6.31
Mexico 22 (2.3%) 3.84±1.45 1.26±0.72 1.95±0.93 1.30±0.51 1.02±0.08 2.84±0.99 2.04
India 21 (2.2%) 5.86±5.73 9.58±5.82 15.71±6.27 1.55±0.43 1.49±0.55 5.29±4.03 6.58
Germany 45 (4.7%) 4.84±4.05 6.92±2.61 9.73±3.90 2.20±1.63 1.38±0.17 5.09±2.00 5.03
China 30 (3.1%) 10.09±6.49 6.93±2.75 10.59±5.73 1.29±0.99 0.83±0.09 5.08±3.75 5.80
Iran 13 (1.4%) 8.25±5.29 6.80±5.72 8.39±7.91 1.53±0.56 1.39±0.53 7.92±6.33 5.71
Greece 18 (1.9%) 3.54±4.18 0.99±0.22 1.15±0.43 1.48±0.09 1.51±0.14 4.09±1.45 2.13
Spain 38 (4.0%) 2.50±1.31 4.83±2.29 6.17±2.48 1.25±0.17 1.25±0.04 4.64±1.45 3.44
Russia 45 (4.7%) 3.61±1.79 3.23±1.01 4.10±2.41 2.18±0.89 2.16±0.69 2.90±0.74 3.03
Others 327 (34.0%) 3.36±2.13 4.31±1.80 5.79±1.31 1.77±0.15 1.57±0.29 3.68±1.00 3.41
ALL 961 (100.0%) 4.29±2.63 5.07±1.92 6.89±1.46 1.85±0.49 1.55±0.16 4.29±1.62 3.99
(a) Performance results evaluated on mAP (%).
Bb-ru mB-c mB-u XRb XRl mT5 Average
Italy 0.3229±0.06 0.3243±0.06 0.3730±0.06 0.2217±0.05 0.2206±0.05 0.3660±0.03 0.3048
U.S. 0.3555±0.05 0.3370±0.07 0.3895±0.03 0.2439±0.06 0.2336±0.03 0.3844±0.03 0.3240
Turkey 0.3129±0.05 0.3006±0.06 0.3554±0.05 0.2018±0.07 0.1915±0.04 0.3472±0.05 0.2849
Japan 0.2609±0.04 0.2622±0.02 0.2906±0.01 0.1619±0.04 0.1618±0.04 0.2712±0.01 0.2348
France 0.3233±0.05 0.2967±0.04 0.3378±0.03 0.2126±0.05 0.2099±0.04 0.3825±0.05 0.2938
U.K. 0.3513±0.08 0.3186±0.07 0.3735±0.03 0.2208±0.08 0.2059±0.05 0.4273±0.07 0.3162
Mexico 0.3264±0.03 0.3010±0.04 0.3348±0.04 0.2480±0.05 0.2431±0.03 0.3815±0.06 0.3058
India 0.2936±0.07 0.3112±0.09 0.4053±0.04 0.1877±0.05 0.1862±0.05 0.3235±0.04 0.2846
Germany 0.3403±0.06 0.3267±0.06 0.3825±0.07 0.2064±0.07 0.1894±0.03 0.3722±0.04 0.3029
China 0.3415±0.08 0.3144±0.04 0.3596±0.04 0.1873±0.06 0.1842±0.05 0.3515±0.04 0.2898
Iran 0.3358±0.07 0.3156±0.05 0.3699±0.06 0.2099±0.06 0.2020±0.04 0.4037±0.05 0.3061
Greece 0.3058±0.11 0.2960±0.05 0.3348±0.02 0.2252±0.04 0.2303±0.05 0.3716±0.02 0.2939
Spain 0.3287±0.04 0.3035±0.06 0.3532±0.03 0.2425±0.04 0.2406±0.03 0.3946±0.04 0.3105
Russia 0.3495±0.03 0.3152±0.05 0.3489±0.04 0.2590±0.05 0.2575±0.05 0.4177±0.03 0.3246
Others 0.3240±0.05 0.3105±0.05 0.3489±0.04 0.2186±0.05 0.2162±0.04 0.3660±0.03 0.2974
ALL 0.3245±0.05 0.3098±0.05 0.3536±0.03 0.2173±0.05 0.2132±0.04 0.3674±0.03 0.2976
(b) Performance results evaluated on mWS.
Table 8: Probing performance comparison with Russian prompts and FmLAMA-ru sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT, “XR” denotes XLM-RoBERTa, “b/l” stands for base/large and “c/u” stands for cased/uncased. Bold and underline represent the best-performing and second-performing cultural group within the same model. The Pearson correlation coefficient between mAP and mWS is 0.70 based on the results above.
Origin Count Bb-ky Bb-kl mB-c mB-u XRb XRl mT5 Avg.
Italy 88 (10.9%) 6.77±2.85 7.01±2.68 1.19±0.15 3.14±2.45 1.90±1.27 2.96±3.44 6.81±1.50 4.25
U.S. 64 (7.9%) 10.72±5.32 10.49±4.97 2.74±1.35 4.84±4.86 3.07±3.62 3.25±3.75 11.65±2.84 6.68
Turkey 19 (2.4%) 3.23±2.85 2.77±0.81 3.58±2.81 4.53±2.97 1.83±0.07 2.37±0.71 1.75±0.39 2.87
Japan 101 (12.5%) 5.56±3.00 5.73±3.43 2.43±1.00 1.94±1.37 3.45±0.68 3.47±0.41 6.77±1.79 4.19
France 67 (8.3%) 2.33±0.74 3.56±1.47 1.40±0.31 2.76±0.17 1.60±0.24 1.62±0.20 2.42±0.12 2.24
U.K. 27 (3.3%) 12.25±6.50 16.39±5.29 4.46±4.18 5.81±5.41 2.11±1.30 3.81±4.73 11.71±4.34 8.08
Mexico 13 (1.6%) 7.40±5.58 4.58±1.71 1.99±0.22 5.58±4.32 1.79±0.17 1.78±0.21 2.28±1.25 3.63
India 32 (4.0%) 4.54±2.64 5.22±1.79 1.20±0.08 2.06±1.10 2.70±3.58 1.68±1.30 5.55±1.09 3.28
Germany 15 (1.9%) 2.26±1.15 3.27±0.71 1.56±0.19 2.74±2.45 2.26±2.37 2.06±1.56 3.51±2.73 2.52
China 54 (6.7%) 6.59±3.41 7.07±3.14 2.10±1.05 2.43±2.01 3.30±3.93 2.80±2.60 6.10±1.59 4.34
Iran 8 (1.0%) 1.74±0.76 8.47±6.64 2.28±0.13 1.50±0.80 1.30±0.48 1.15±0.21 1.56±0.53 2.57
Greece 9 (1.1%) 1.27±0.27 6.00±3.98 2.06±0.09 2.31±0.56 0.97±0.15 1.12±0.11 5.67±4.15 2.77
Spain 21 (2.6%) 2.53±1.12 2.95±1.22 3.84±2.58 2.53±2.12 3.04±2.33 3.87±3.90 4.26±2.11 3.29
Russia 8 (1.0%) 2.04±1.88 3.98±2.35 1.28±0.28 1.89±0.78 3.05±4.75 1.71±1.19 1.82±0.33 2.25
Others 281 (34.8%) 4.46±1.86 5.12±1.67 1.84±0.31 1.93±0.99 1.87±1.11 2.00±1.21 4.24±0.63 3.07
ALL 807 (100.0%) 5.42±2.33 6.09±2.16 2.05±0.71 2.68±1.74 2.31±1.52 2.49±1.70 5.56±0.81 3.80
Table 9: Probing performance comparison on mAP (%) with Korean prompts and FmLAMA-ko sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT, “XR” denotes XLM-RoBERTa, “b/l” stands for base/large and “c/u” stands for cased/uncased. Bold and underline represent the best-performing and second-performing cultural groups within the same model.
Origin Count Bb-zh mB-c mB-u XRb XRl mT5 Avg.
Italy 49 (5.8%) 13.30±4.45 10.57±2.58 14.37±2.34 10.52±3.21 11.61±3.41 10.15±1.19 11.75
U.S. 101 (11.9%) 15.92±3.83 14.10±2.67 16.08±1.70 9.00±5.25 9.31±4.08 12.67±0.77 12.85
Turkey 12 (1.4%) 8.26±1.41 4.10±1.61 4.91±1.53 2.43±0.85 2.02±0.54 4.94±0.48 4.44
Japan 114 (13.4%) 11.51±4.77 12.29±2.43 12.89±2.23 7.31±4.70 8.32±4.04 7.85±0.48 10.03
France 70 (8.2%) 15.34±4.47 15.31±3.04 15.29±2.19 10.36±5.22 12.79±3.97 7.90±0.97 12.83
U.K. 38 (4.5%) 23.26±5.66 15.85±4.90 18.49±4.34 14.03±6.94 14.95±4.25 9.00±2.26 15.93
Mexico 19 (2.2%) 17.20±3.14 12.52±1.39 12.63±1.61 10.68±3.44 13.99±1.16 15.19±1.90 13.70
India 24 (2.8%) 5.69±2.14 4.96±2.12 6.79±2.52 3.50±2.36 5.27±1.84 1.01±0.11 4.54
Germany 16 (1.9%) 6.97±2.09 11.28±1.30 13.08±3.16 5.61±4.45 5.74±2.34 7.55±0.64 8.37
China 87 (10.2%) 16.68±7.05 15.47±4.87 17.70±6.07 9.20±5.50 12.71±6.72 8.33±1.34 13.35
Iran 7 (0.8%) 9.25±1.04 5.60±1.05 11.30±6.38 7.38±6.42 8.71±6.74 3.21±0.09 7.58
Greece 7 (0.8%) 4.29±0.79 2.13±0.70 1.93±0.28 2.70±1.17 5.58±2.37 4.35±0.48 3.50
Spain 12 (1.4%) 11.98±1.11 7.52±2.42 10.18±1.07 7.55±2.97 9.46±3.33 11.65±1.92 9.72
Russia 8 (0.9%) 5.48±0.98 2.82±0.84 3.79±0.69 2.56±0.98 3.65±1.62 2.11±0.43 3.40
Others 251 (30.8%) 10.55±3.73 9.90±2.97 10.28±1.60 6.59±3.07 9.38±3.56 5.70±0.41 8.73
ALL 815 (100.0%) 12.99±3.96 11.78±2.65 13.01±2.06 8.05±4.00 9.98±3.69 7.88±0.55 10.62
Table 10: Probing performance comparison with Chinese prompts and FmLAMA-zh sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT, “XR” denotes XLM-RoBERTa, “b/l” stands for base/large and “c/u” stands for cased/uncased. Bold and underline represent the best-performing and second-performing cultural groups within the same model.
Origin mB-c mB-u mT5 Average
Italy 9.70±4.24 10.42±6.42 12.61±2.28 10.91
U.S. 14.91±4.71 12.86±5.63 11.45±3.22 13.07
Japan 8.27±3.62 9.04±5.07 6.99±3.13 8.10
France 8.26±3.15 8.13±4.40 5.00±2.18 7.13
U.K. 19.84±7.17 13.97±9.13 13.11±7.97 15.64
India 13.87±4.46 13.94±7.00 8.84±3.44 12.22
China 14.89±3.74 13.23±5.08 9.04±2.73 12.39
Indonesia 8.33±3.68 6.75±2.38 6.79±2.12 7.29
ALL 11.97±3.27 10.98±4.65 9.03±2.77 10.66
(a) EN_en: mAP (%)
Origin mB-c mB-u mT5 Average
Italy 5.77±3.04 6.18±3.57 4.21±1.12 5.39
U.S. 5.11±2.31 5.75±1.41 2.62±0.63 4.49
Japan 10.51±2.87 9.66±4.55 2.75±1.23 7.64
France 5.02±2.41 6.06±1.66 2.76±0.54 4.61
U.K. 3.14±1.06 3.77±0.98 2.93±0.54 3.28
India 14.43±9.56 16.22±7.65 4.05±1.59 11.57
China 3.87±1.39 6.08±3.77 2.85±1.04 4.27
Indonesia 10.60±1.80 8.76±3.04 3.32±0.90 7.56
ALL 6.77±2.26 7.28±1.95 2.99±0.65 5.68
(b) EN_zh: mAP (%)
Origin mB-c mB-u mT5 Average
Italy 7.25±3.72 6.66±5.45 7.15±1.16 7.02
U.S. 13.23±3.64 7.69±5.49 11.42±2.11 10.78
Japan 12.19±3.70 6.45±4.36 7.61±1.81 8.75
France 11.64±4.46 7.25±5.96 7.66±2.84 8.85
U.K. 13.60±6.01 9.17±7.54 11.34±4.36 11.37
India 3.70±1.51 3.03±1.66 1.57±0.18 2.77
China 15.04±3.69 9.09±8.25 8.80±3.81 10.98
Indonesia 9.65±1.93 4.88±2.55 5.41±2.06 6.65
ALL 11.88±3.49 7.22±5.48 8.40±2.22 9.17
(c) ZH_zh: mAP (%)
Origin mB-c mB-u mT5 Average
Italy 1.53±0.20 2.28±0.25 2.52±0.56 2.11
U.S. 2.02±0.21 3.46±0.49 3.45±0.69 2.98
Japan 2.59±0.44 2.74±0.73 2.65±0.15 2.66
France 2.04±0.60 2.41±0.58 2.42±0.44 2.29
U.K. 1.42±0.22 3.40±1.27 2.29±0.35 2.37
India 2.03±0.37 2.53±0.77 2.23±0.91 2.26
China 2.85±0.44 2.99±1.34 2.29±0.56 2.71
Indonesia 2.04±0.27 2.36±0.36 1.25±0.20 1.88
ALL 2.20±0.33 2.85±0.64 2.58±0.34 2.54
(d) ZH_en: mAP (%)
Table 11: Code-switching analysis: Probing results by using prompts without introducing cultural background.
Origin mB-c mB-u mT5 Average
Italy 9.20±4.55 9.97±4.71 11.27±2.42 10.15
U.S. 15.49±4.37 13.05±5.40 11.37±4.10 13.30
Japan 11.81±3.16 11.84±3.99 10.11±3.43 11.25
France 8.29±2.81 8.93±4.29 5.92±1.38 7.71
U.K. 20.02±9.00 15.81±7.65 15.79±6.17 17.21
India 17.95±2.38 21.88±7.42 11.51±4.65 17.11
China 14.77±3.49 13.75±6.35 9.13±3.70 12.55
Indonesia 9.02±1.01 8.75±2.36 8.11±3.19 8.63
ALL 13.06±3.26 12.42±4.46 10.12±3.16 11.87
(a) EN_en: mAP (%)
Origin mB-c mB-u mT5 Average
Italy 7.57±1.67 6.79±1.71 4.18±1.02 6.18
U.S. 6.01±1.71 7.51±1.08 2.86±0.55 5.46
Japan 13.96±2.05 10.45±3.31 4.77±1.61 9.73
France 5.28±1.72 6.63±0.78 2.92±0.70 4.94
U.K. 3.19±0.91 3.60±0.77 4.44±1.47 3.74
India 23.58±2.64 19.51±6.39 6.59±2.59 16.56
China 5.23±0.14 6.39±3.55 3.98±1.08 5.20
Indonesia 10.12±1.01 7.91±2.69 6.35±1.81 8.13
ALL 8.57±1.25 8.09±1.35 4.10±0.98 6.92
(b) EN_zh: mAP (%)
Origin mB-c mB-u mT5 Average
Italy 6.35±3.08 7.67±4.89 6.33±1.80 6.78
U.S. 13.60±3.06 8.50±5.41 10.17±1.91 10.76
Japan 13.02±2.43 7.32±4.47 8.83±1.60 9.72
France 12.12±4.46 9.00±5.67 7.48±2.82 9.53
U.K. 13.97±6.02 8.21±7.81 10.87±4.47 11.02
India 4.60±1.88 3.38±2.06 1.82±0.30 3.27
China 20.49±4.29 10.22±8.87 9.71±4.53 13.47
Indonesia 9.09±3.50 4.12±2.30 5.53±1.74 6.25
ALL 13.07±3.13 7.99±5.34 8.46±2.24 9.84
(c) ZH_zh: mAP (%)
Origin mB-c mB-u mT5 Average
Italy 1.57±0.18 2.73±0.63 2.77±0.61 2.36
U.S. 2.29±0.13 3.61±1.16 4.30±0.73 3.40
Japan 2.79±0.65 2.69±1.35 2.35±0.60 2.61
France 3.02±0.86 2.56±0.70 2.25±0.20 2.61
U.K. 1.77±0.21 2.38±1.43 2.24±0.38 2.13
India 2.30±0.51 2.85±1.29 2.30±0.86 2.48
China 3.12±0.71 2.39±1.70 2.31±0.45 2.61
Indonesia 2.39±0.46 2.28±0.77 1.20±0.12 1.96
ALL 2.54±0.42 2.77±1.07 2.68±0.21 2.66
(d) ZH_en: mAP (%)
Table 12: Code-switching analysis: Probing results by using prompts with introducing cultural background.
Origin Llama2-7b-chat-hf Llama2-13b-chat-hf
EN_en EN_zh ZH_zh ZH_en EN_en EN_zh ZH_zh ZH_en
Italy 42.86 20.41 2.04 2.04 48.98 30.61 6.12 10.20
U.S. 37.62 13.86 4.95 4.95 49.50 27.72 5.94 12.87
Japan 14.91 16.67 3.51 4.39 20.18 15.79 4.39 6.14
France 39.71 22.06 5.88 11.76 48.53 26.47 2.94 10.29
U.K. 36.84 18.42 15.79 5.26 39.47 21.05 15.79 7.89
India 20.83 12.50 4.17 4.17 33.33 12.50 12.50 16.67
China 17.65 23.53 10.59 15.29 21.18 24.71 16.47 15.29
Indonesia 6.90 10.34 0 6.90 24.14 13.79 3.45 3.45
All 27.36 17.91 5.91 7.28 35.04 22.64 7.87 10.43
(a) Performance results of Llama2-chat.
Origin vicuna-7b vicuna-13b
EN_en EN_zh ZH_zh ZH_en EN_en EN_zh ZH_zh ZH_en
Italy 51.02 26.53 8.16 12.24 42.86 30.61 18.37 22.45
U.S. 49.50 27.72 17.82 13.86 37.62 25.74 20.79 22.77
Japan 21.93 14.04 7.02 7.02 16.67 14.04 9.65 7.02
France 44.12 23.53 11.76 13.24 39.71 20.59 14.71 17.65
U.K. 47.37 21.05 15.79 10.53 34.21 31.58 23.68 26.32
India 41.67 4.17 12.50 12.50 16.67 16.67 12.50 12.50
China 23.53 18.82 27.06 15.29 17.65 18.82 20.00 22.35
Indonesia 20.69 6.90 10.34 13.79 20.69 10.34 13.79 13.79
All 36.22 19.69 14.37 12.01 28.15 20.87 16.54 17.72
(b) Performance results of vicuna.
Origin gpt-3.5-turbo
EN_en EN_zh ZH_zh ZH_en
Italy 69.39 48.98 20.41 26.53
U.S. 60.40 39.60 22.77 25.74
Japan 28.95 31.58 10.53 16.67
France 63.24 33.82 17.65 27.94
U.K. 57.89 36.84 26.32 26.32
India 54.17 12.50 12.50 16.67
China 30.59 34.12 31.76 37.65
Indonesia 34.48 13.79 20.69 13.79
All 47.64 34.06 20.28 25.00
(c) Performance results of gpt-3.5-turbo.
Table 13: Accuracy (%) of decoder-only LLMs’ probing results in each cultural group.