### v3

# Does Mapo Tofu Contain Coffee? Probing LLMs for Food-related Cultural Knowledge

Abstract
Recent studies have highlighted the presence of cultural biases in Large Language Models (LLMs), yet often lack a robust methodology to dissect these phenomena comprehensively. Our work aims to bridge this gap by delving into the Food domain—a universally relevant yet culturally diverse aspect of human life. We introduce FmLAMA, a multilingual dataset centered on food-related cultural facts and variations in food practices. We analyze LLMs across various architectures and configurations, evaluating their performance in both monolingual and multilingual settings. By leveraging templates in six different languages, we investigate how LLMs interact with language-specific and cultural knowledge. Our findings reveal that (1) LLMs demonstrate a pronounced bias towards food knowledge prevalent in the United States; (2) Incorporating relevant cultural context significantly improves LLMs’ ability to access cultural knowledge; (3) The efficacy of LLMs in capturing cultural nuances is highly dependent on the interplay between the probing language, the specific model architecture, and the cultural context in question. This research underscores the complexity of integrating cultural understanding into LLMs and emphasizes the importance of culturally diverse datasets to mitigate biases and enhance model performance across different cultural domains.
Does Mapo Tofu Contain Coffee?
Probing LLMs for Food-related Cultural Knowledge
Li Zhou1,2, Taelin Karidi3, Wanlong Liu4, Nicolas Garneau5, Yong Cao6, Wenyu Chen4, Haizhou Li1,2, Daniel Hershcovich5 1The Chinese University of Hong Kong, Shenzhen 2Shenzhen Research Institute of Big Data 3Hebrew University of Jerusalem 4University of Electronic Science and Technology of China, 5University of Copenhagen, 6University of Tübingen, Tübingen AI Center {lizhou21, haizhouli}@cuhk.edu.cn, taelin.karidi@mail.huji.ac.il, liuwanlong@std.uestc.edu.cn, {nicolas.garneau, dh}@di.ku.dk, yongcao2018@gmail.com, cwy@uestc.edu.cn
1 Introduction
Figure 1: Summary of the various aspects of our work.
[…]
Trained on vast datasets, Large Language Models (LLMs) encode a wide array of knowledge, but also face challenges with various biases, such as those related to gender Savoldi et al. (2021); Kaneko et al. (2022), belief Søgaard (2021); González et al. (2021); Lent and Søgaard (2021), and culture Cao et al. (2023); Deshpande et al. (2022); Yin et al. (2022); Mukherjee et al. (2024); Singh et al. (2024). In this paper, we adopt food-domain knowledge data as cultural semantic proxy, “country of origin” as cultural demographic proxy Adilazuarda et al. (2024), and address the following research questions: (1) What cultural biases exist within the cultural knowledge contained in LLMs? (2) How can we best prompt and evaluate LLMs to elicit culture-specific knowledge? Specifically, we first design an automated construction method for the dataset FmLAMA3). Next, we introduce the cultural knowledge probing method (§4), encompassing the probing task, template and metric. We then conduct experiments (§5) and analysis (§6) to address our proposed research questions. Finally, we perform an error analysis on the probing results (§7). Figure 1 provides an overview of our work’s various aspects, and our contributions are:
  • We present FmLAMA, a pioneering dataset focused on the food domain, which is inherently rich in cultural diversity. This dataset is a multifaceted tool for probing LLMs across cultures and languages.
  • We propose novel metrics aimed at evaluating LLMs’ capacity to accurately and sensitively probe cultural knowledge, utilizing a combination of automatic and manual evaluation methods.
  • We analyze the impact of integrating cultural context and language specificity in prompts, offering insights to optimize LLMs for equitable cross-cultural knowledge retrieval.
Our methodology for automated collection of cultural knowledge corpora extends the analysis potential in other domains, broadening the scope of research on cultural biases in LLMs.  [^1] Data and code are available in https://github.com/lizhou21/FmLAMA-master.
2 Related Work
Cultural knowledge datasets.
Cultural knowledge, encompassing the customs, beliefs, traditions, and practices of a culture, is crucial yet challenging to encapsulate. While some researchers focus on manually curating cultural knowledge datasets, others evaluate LLMs’ performance on culturally related tasks. Yin et al. (2022) and Palta and Rudinger (2023) have developed benchmarks such as geo-diverse prompts and food-custom datasets (FORK) to probe cultural biases in commonsense reasoning systems. However, manual dataset construction is inefficient and hard to scale, prompting a shift towards automated methods. For instance, StereoKG Deshpande et al. (2022) offers a scalable knowledge graph that blends cultural knowledge with stereotypes, and Candle Nguyen et al. (2023) extracts cultural commonsense knowledge from the web, organizing it into clusters. Despite these advances, the variability in data representation—from sentences to triplets using OpenIE—poses challenges for consistency and noise control in knowledge probing. Keleg and Magdy (2023) aims to mitigate this by selecting culturally diverse factual triples from Wikidata, focusing mainly on explicit country information. In contrast, our work proposes an automated, efficient approach to constructing a cultural knowledge dataset in a uniform triplet format, addressing the limitations of existing methods and focusing on implicit cultural knowledge. Appendix A.1 shows a comparison of our dataset with existing ones.
Knowledge probing.
Deciphering the knowledge encoded by LLMs poses significant challenges due to their opaque nature , early benchmarks like LAMA Petroni et al. (2019) sought to quantify the factual knowledge in English LLMs, while ParaRel Elazar et al. (2021) highlighted their consistency issues. Subsequent efforts as mLAMA Kassner et al. (2021) and mParaRel Fierro and Søgaard (2022) expanded these benchmarks multilingually, though such methods often focus on single-word entities, limiting their depth of assessment. To address these shortcomings, newer studies Shin et al. (2020); Zhong et al. (2021); Meng et al. (2022) have evolved towards eliciting more comprehensive factual knowledge, including multi-word entities, with Jiang et al. (2020a) developing algorithms for multi-token predictions. LPAQA Jiang et al. (2020b) further refines this by optimizing prompt discovery for more accurate knowledge probing. Our work builds on this foundation, targeting multi-token probing within the food domain, characterized by complex expressions like Trigonella foenum-graecum.
3 FmLAMA Construction
To assess whether LLMs encode and access cultural information, we develop FmLAMA, a multicultural, multilingual dataset focusing on culinary knowledge. The designed framework can be adapted to other cultural domains.
Step #1: Obtain countries set.
Following Zhou et al. (2023), we use countries of food origin to delineate cultural groups. This method leverages countries as proxies for cultural identity, encapsulating diverse traditions, values, and norms that reflect the breadth of human civilizations across geographical boundaries Minkov and Hofstede (2012); Peterson et al. (2018).
Step #2: Acquire food instances.
We utilize SPARQL to query Wikidata, extracting a vast array of food-related data. This approach exploits Wikidata’s RDF triple structure to gather detailed information on food instances, offering a rich source of comprehensive food knowledge.
i. Class. For our food-focused dataset, we concentrate on the dish class and employ two approaches to find food instances:
  • Explicit instance of dish, e.g., bouillabaisse.
[…]
This enables comprehensive inclusion of food instances, represented as $I\xrightarrow{(\text{instance of}|\text{subclass of})+}$ dish, where ‘||||’ denotes “or”, and ‘$+$’ is “one or more”.
ii. Cultural group. We organize food instances by their origin, applying these strategies:
[…]
We exclude dishes with multiple origin countries to maintain cultural specificity.
iii. Properties included. The property “has part(s)” identifies food ingredients for a dish. Additional properties like “made from material” and “image” are collected to support future research (e.g., multimodal), though they are not utilized in this study. In each instance, the descriptive language for all attributes remains consistent.
Ultimately, our constructed dataset, FmLAMA, comprises 33,600 dish instances, detailed by name, origin, ingredients, and optionally, materials and images, encompassing 128 cultural country groups and 250 languages. The average number of ingredients in FmLAMA dataset is 2.04. Examples and statistics are provided in Appendix A.
4 Cultural Knowledge Probing
4.1 Probing task
We adopt Word Prediction (WP) as the knowledge-probing task, following previous work Fierro and Søgaard (2022); Wu et al. (2023); Fierro et al. (2024b). Firstly, we manually design a prompt template $t$ focused on the core attribute “has part(s)”, illustrating a connection between a dish (subject) and its ingredient(s) (object). WP is usually implemented as a candidate retrieval problem. The candidate set consists of all objects in the focused, filtered sub-dataset of FmLAMA. [^2] As the size of the filtered sub-dataset increases, the candidate object set also expands, leading to greater difficulty in probing. Consequently, in this paper, results obtained by probing across different filtered sub-datasets cannot be used for horizontal comparison. The primary objective is to utilize LLMs to obtain the probability of each candidate $C$ and subsequently rank the predicted objects based on these probabilities.
MASK operation.
Using subject-object tuples ([X], [Y]) as queries, we probe LLMs by replacing the subject and masking the object. Considering that each candidate object is tokenized into $d$ subtokens $\left\{c_{1},\cdots,c_{d}\right\}$ by LLMs correspondingly, we apply [MASK] token of varying lengths to the objects within each query. $D$ queries are constructed for each dish subject $X$ based on the same template $t$, $D$ is the maximum number of object tokens, each query $Q_{d}^{X}$ is defined as:
$Q_{d}^{X}=t\left(X,\left[\mathrm{MASK}\right]*d\right),d\in\left[1,D\right]\vspace{-3pt}.$ (1)
Probability acquisition.
We use Mean Pooling [^3] Mean Pooling has been shown to be superior to other pooling methods for multi-token probing Wu et al. (2023). method to obtain the prediction probability of each candidate. Specifically, for candidate object $C=\left\{c_{1},\cdots,c_{d}\right\}$ of length $d$, its probability is obtained from the likelihoods associated with the [MASK] tokens in $Q_{d}^{X}$, and the predicted probability of $C$ is calculated as the average of the probabilities of composing its subtokens:
$P\left(Q_{d}^{X},C\right)=\frac{1}{d}\sum_{i=1}^{d}{p\left(Q_{d}^{X},\left[\mathrm{MASK}\right]_{i}=c_{i}\right)}\vspace{-3pt},$ (2)
where $p\left(\cdot\right)$ is obtained after the softmax operation.
4.2 Probing template
Considering LLMs produce varied predictions based on prompt framing Elazar et al. (2021); Wang et al. (2023), we craft five templates conveying identical meanings in each prompt language. For example, “[X] is a dish made with [Y]”, “[X] is a type of food that includes [Y]”.
Cultural information.
To explore the impact of introducing explicit cultural context on LLMs’ ability to access cultural knowledge, we enhance the basic templates by integrating country information. For instance, “In [C], [X] is a dish made with [Y]”, where [C] denotes the country of origin for the dish [X], [Y] indicates the ingredient object.
Multilingual prompts.
To explore how different prompt language settings affect LLMs’ cultural knowledge probing abilities, we craft prompts in six languages written by native speakers, including English (en), Chinese (zh), Arabic (ar), Korean (ko), Russian (ru), and Hebrew (he). These languages span 4 different language families – Indo-European (English, Russian), Semitic (Hebrew, Arabic), Altaic (Korean), and Sino-Tibetian (Chinese), and are spoken by more than 2.356 billion speakers. Furthermore, these languages represent cultural diversity, being spoken on different continents by groups with rich and distinct cultural backgrounds. These prompt illustrations and language details are depicted in Appendix B.
Code-switching.
[…]
4.3 Probing Metric
[…]
Given these constraints, we introduce two automatic metrics: the absolute-match metric, Mean Average Precision (mAP), and the fuzzy-match metric, Mean Word Similarity (mWS), along with a Manual Evaluation Score (MES) that incorporates both LLM-simulated and real human assessments for implementation. All mAP, mWS, and MES metrics are in the range $[0,1]$.
Mean Average Precision.
mAP is widely used in information retrieval settings, assessing the relevance of predicted objects (in our case ingredients) only when they precisely match the reference ones. The precision at rank $k$ (P@k) for a given food instance $i$ is defined as follows:
$\mathrm{P@k}={\frac{\left|\mathrm{ing}_{i}\cap\mathrm{topk}_{i}\right|}{k}}\vspace{-3pt},$ (3)
where $\mathrm{ing}_{i}$ is the reference ingredients set, and $\mathrm{topk}_{i}$ signifies the set of top-$k$ objects with the highest predicted probability of belonging to food item $i$ by LLMs. Then the average precision of food item $i$ is computed as follows:
$\mathrm{AP}_{i}=\frac{1}{\left|\mathrm{ing}_{i}\right|}\sum_{k=1}^{n}{\mathrm{P}@\mathrm{k}\times\mathrm{rel}@\mathrm{k}}\vspace{-3pt},$ (4)
where $n$ refers to the size of the candidate object set, and $\mathrm{rel}@\mathrm{k}$ is 1 if the object at rank $k$ is relevant to food item $i$, otherwise 0. Finally, we compute the mAP in the following way:
$\mathrm{mAP}=\frac{1}{\left|G\right|}\sum_{i\in G}{\mathrm{AP}_{i}}\vspace{-3pt},$ (5)
where $G$ represents a dish group we are focusing on (i.e. a subset of FmLAMA).
Mean Word Similarity.
mWS is defined based on the semantic similarity between predicted and reference objects. First, we define the similarity score $S\left(i,g\right)$ for each ingredient $g$ within each dish $i$. Only the predicted objects in the top-$l$ rankings of the model prediction that are most similar to $g$ contribute to the evaluation score, where $l=\left|\mathrm{ing}_{i}\right|$. $S\left(i,g\right)$ is defined as follows:
$S@l\left(i,g\right)=\underset{p\in\mathrm{topl}_{i}}{\max}\left[\cos\left(w_{g},w_{p}\right)\right],g\in\mathrm{ing}_{i}\vspace{-3pt},$ (6)
where $w_{g}$ and $w_{p}$ are the Fasttext Bojanowski et al. (2017); Joulin et al. (2016) vectors for the ingredient $g$ and the predicted object $p$, respectively, and $\cos\left(\cdot\right)$ is the cosine similarity function. Then we can compute the probing similarity $WS_{i}$ for each food instance $i$ and mWS for the targeted food group as follows:
$\mathrm{WS}_{i}$ $=$ $\frac{1}{\left|\mathrm{ing}_{i}\right|}\sum_{g\in\mathrm{ing}_{i}}{S@l\left(i,g\right)}$ (7)
$\mathrm{mWS}$ $=$ $\frac{1}{\left|G\right|}\sum_{i\in G}{\mathrm{WS}_{i}}$ (8)
Encoder-only LLMs Encoder-Decoder LLMs Decoder-only LLMs
Origin Count Dumb Base Bb Bl mB T5 mT5 Qwen2 Llama2 Llama3 Avg.
Italy 215 (13.9%) 18.14 8.12±1.14 7.52±0.43 9.31±1.33 4.54±0.56 6.73±1.34 26.82±4.36 19.37±3.68 29.62±3.30 14.00
U.S. 285 (18.4%) 11.10 19.98±2.00 18.83±1.95 19.91±5.88 10.24±2.62 16.50±4.86 32.21±2.05 21.64±5.75 36.58±3.58 21.99
Turkey 98 (6.3%) 12.90 11.80±1.03 9.84±3.81 14.98±3.89 6.10±1.57 10.25±2.91 25.22±1.46 22.96±4.04 30.01±2.30 16.39
Japan 186 (12.0%) 9.35 11.10±1.19 9.14±1.81 11.52±1.89 5.45±1.90 6.38±1.00 23.15±2.01 20.24±3.10 25.47±1.73 14.06
France 175 (11.3%) 16.50 9.33±1.24 7.51±0.90 8.80±2.34 3.86±1.62 4.57±1.11 29.56±2.38 21.95±4.83 30.01±2.19 14.45
U.K. 83 (5.4%) 18.67 12.47±1.77 13.72±2.20 14.20±4.01 6.16±2.89 9.88±3.06 29.06±2.47 21.82±3.43 32.15±4.06 17.43
Mexico 57 (3.7%) 9.40 8.25±1.04 10.35±2.88 8.87±2.18 3.41±0.49 5.96±2.36 25.14±1.58 22.03±5.65 28.97±3.54 14.12
India 132 (8.5%) 11.57 18.42±1.35 16.94±3.10 17.51±1.75 9.18±2.98 7.80±1.15 35.82±1.79 33.88±5.00 37.62±1.89 22.15
Germany 57 (3.7%) 14.02 8.54±1.77 7.84±1.01 9.00±2.39 4.38±1.48 6.68±1.56 29.05±2.63 20.42±5.56 28.86±3.78 14.35
China 97 (6.3%) 8.60 15.06±2.45 14.94±1.37 15.24±5.31 7.80±1.97 12.39±2.62 28.96±3.15 21.02±4.78 30.39±2.70 18.23
Iran 21 (1.4%) 12.60 9.16±1.75 8.19±1.51 13.56±3.16 9.00±2.04 6.21±0.15 40.26±8.19 35.36±9.81 45.72±3.65 20.93
Greece 21 (1.4%) 15.07 4.09±1.10 3.48±1.45 4.85±1.37 3.81±1.09 1.32±0.15 31.68±8.85 15.88±2.54 31.60±3.17 12.09
Spain 95 (6.1%) 16.05 9.18±0.93 7.18±1.04 7.42±1.51 3.98±1.37 3.33±0.46 25.21±5.07 17.00±4.23 25.36±1.68 12.33
Russia 27 (1.7%) 10.64 4.72±0.98 7.10±2.81 6.44±1.97 1.41±0.60 2.21±0.33 14.83±4.80 10.47±0.98 11.77±0.59 7.37
ALL 1549 (100.0%) 13.29 12.56±1.06 11.58±1.29 12.96±2.84 6.32±1.42 8.68±2.03 28.56±2.27 21.89±4.06 30.96±2.09 16.69
Coefficient of Variation (CV) 24.15 40.87 40.76 36.89 43.29 54.29 20.39 28.27 23.84 -
(a) Performance results evaluated using mAP (%). CV indicates the extent of cultural bias in each model’s performance.
Origin Encoder-only LLMs Encoder-Decoder LLMs Decoder-only LLMs Avg.
Bb Bl mB T5 mT5 Qwen2 Llama2 Llama3
Italy 0.3813±0.02 0.3624±0.01 0.3573±0.02 0.3129±0.02 0.3195±0.01 0.4930±0.02 0.4510±0.02 0.5057±0.01 0.3979
U.S. 0.4461±0.02 0.4358±0.02 0.4219±0.06 0.3516±0.02 0.4076±0.03 0.5147±0.02 0.4656±0.04 0.5411±0.02 0.4481
Turkey 0.4130±0.02 0.3886±0.03 0.3954±0.04 0.3387±0.03 0.3419±0.02 0.4689±0.01 0.4612±0.03 0.5022±0.01 0.4137
Japan 0.3778±0.02 0.3501±0.03 0.3446±0.03 0.3059±0.03 0.2932±0.01 0.4416±0.01 0.4304±0.02 0.4436±0.01 0.3734
France 0.3750±0.02 0.3660±0.02 0.3497±0.04 0.3279±0.02 0.3367±0.01 0.5201±0.02 0.4697±0.03 0.5099±0.01 0.4069
U.K. 0.3708±0.02 0.3775±0.02 0.3583±0.04 0.3031±0.02 0.3320±0.03 0.4532±0.03 0.4291±0.01 0.4620±0.04 0.3857
Mexico 0.3509±0.01 0.3654±0.03 0.3255±0.03 0.3287±0.01 0.3262±0.02 0.4547±0.02 0.4762±0.04 0.4863±0.03 0.3892
India 0.4402±0.01 0.4193±0.04 0.4135±0.02 0.3380±0.04 0.3141±0.01 0.5106±0.00 0.5140±0.03 0.5190±0.02 0.4336
Germany 0.3699±0.03 0.3492±0.03 0.3698±0.03 0.2987±0.02 0.3358±0.01 0.4987±0.02 0.4758±0.05 0.4983±0.01 0.3995
China 0.3676±0.02 0.3532±0.02 0.3317±0.05 0.3368±0.03 0.3670±0.03 0.4521±0.04 0.4285±0.02 0.4671±0.02 0.3880
Iran 0.3950±0.03 0.3747±0.03 0.3972±0.03 0.3384±0.04 0.3303±0.02 0.5510±0.06 0.5498±0.07 0.6086±0.03 0.4431
Greece 0.3854±0.03 0.3707±0.03 0.3461±0.02 0.3569±0.04 0.3257±0.01 0.5415±0.08 0.4459±0.02 0.5527±0.02 0.4156
Spain 0.3761±0.01 0.3386±0.02 0.3290±0.02 0.3215±0.02 0.2923±0.01 0.4630±0.03 0.4247±0.03 0.4641±0.01 0.3762
Russia 0.3686±0.01 0.3778±0.03 0.3286±0.01 0.3386±0.01 0.2912±0.02 0.4272±0.04 0.3969±0.02 0.4053±0.00 0.3668
ALL 0.3959±0.01 0.3803±0.02 0.3699±0.03 0.3276±0.02 0.3380±0.01 0.4864±0.01 0.4573±0.03 0.4981±0.01 0.4067
Corr. 0.35 0.59 0.38 0.57 0.83 0.86 0.71 0.92 -
(b) Performance results evaluated using mWS with Fasttext. Corr. represents the correlation between the FastText and BERT base embedding evaluation results.
Table 1: Automatic Evaluation: Probing performance comparison with English prompts and FmLAMA-en sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT, and “b/l” stands for base/large. Bold and underline indicate the best and second-best performing cultural groups, respectively, for each model (within each column). The average Pearson correlation between mAP and mWS across LLMs is 0.72.
Bb Bl mB T5
mAP 0.60 0.49 0.48 0.56
mWS 0.48 0.38 0.36 0.41
mT5 Qwen2 Llama2 Llama3
mAP 0.38 0.05 0.04 0.13
mWS -0.06 0.05 -0.04 0.02
Table 2: Correlation between Country Dish Counts and Probing Performance.
Origin GPT-4o Human Origin GPT-4o Human Origin GPT-4o Human
Italy 52.40 55.70 U.K. 64.12 72.33 Iran 53.39 51.70
U.S. 70.06 70.26 Mexico 47.05 51.95 Greece 47.68 51.23
Turkey 55.19 53.59 India 64.22 63.72 Spain 43.02 53.16
Japan 50.76 56.07 Germany 51.71 59.02 Russia 31.42 44.86
France 52.03 58.28 China 61.52 68.96 All 56.58 -
Table 3: Manual Evaluation: Llama3 probing results evaluated using MES. The human evaluation score is the average MES from three evaluators per country group. Since evaluators vary by country group, the overall result (ALL) is shown as “-”.
Manual Evaluation Score.
MES considers only the top-$l$ predicted ingredients for evaluation in each dish, similar to the mWS metric, where $l$ is the number of reference ingredients. A predicted ingredient is considered correct if it satisfies any of the following criteria: (1) Direct Match: The ingredient is exactly in the Wikidata reference; (2) Substitutability: The ingredient can replace one in the reference; (3) Missing Traditional Ingredient: The ingredient is traditionally or commonly used in the dish, but not listed in the reference. Given a manual evaluator $\mathrm{M}\left(\cdot\right)$, the MES is defined as follows:
$\mathrm{S}_{i}$ $=$ $\frac{1}{l}\sum_{p\in\mathrm{topl}_{i}}{\mathrm{M}\left(p,\mathrm{ing}_{i}\right)}$ (9)
$\mathrm{MES}$ $=$ $\frac{1}{\left|G\right|}\sum_{i\in\mathrm{G}}{\mathrm{S}_{i}}$ (10)
where $\mathrm{M}\left(\cdot\right)$ is 1 if the top-$l$ predicted ingredient $p$ is evaluated as meeting the criteria; otherwise, 0.
5 Experiments: Existing Cultural Biases
5.1 Experimental setup
Baselines
We explore encoder-only LLMs, including BERT Devlin et al. (2019) and mBERT, encoder-decoder LLMs, such as T5 Raffel et al. (2020) and mT5 Xue et al. (2020), as well as decoder-only LLMs like Qwen2 Yang et al. (2024), Llama2 Touvron et al. (2023), and Llama3 AI@Meta (2024). Of these, BERT and T5 are English monolingual LLMs, while the remaining five are multilingual LLMs.  [^4] Monolingual LLM configurations, as well as results for the five languages on their respective filtered sub-datasets, are provided in the Appendix D.1. Additionally, we employ a ‘dumb’ baseline, in which the model is assumed to consistently predict the top-10 most common ingredients for each dish.
Cultural groups
We identify the top 30 countries with the most dishes across six languages, take their intersection to ensure sufficient data, consider geographical diversity, and ultimately narrow our focus to 14 cultural groups, as shown in Table 1.
Metrics
We evaluate all LLMs using two automated metrics. Due to cost constraints, we conduct manual evaluation only on the LLM that achieve the highest probing performance in the automated metrics. Specifically, for the evaluator $\mathrm{M}\left(\cdot\right)$ used to compute MES, we utilize GPT-4o OpenAI et al. (2024) [^5] Version: gpt-4o-2024-08-06 as a simulated evaluator and recruit real human evaluators from Prolific. [^6] https://www.prolific.co/ For each country group, we hire 3 evaluators familiar with the related cuisine, with a pay rate of £9/hour, considered mid-level on the platform. Details on the GPT-4o evaluation prompts and the human evaluation platform are provided in the Appendix C.
5.2 Automatic evaluation results
The probing results based on English prompt on the filtered dataset FmLAMA-en are illustrated in Table 1. The average Pearson correlation between mAP and mWS across all LLMs is 0.72, indicating a strong positive relationship and demonstrating that the experimental results are consistent across both automatic metrics. Overall, without fine-tuning, decoder-only LLMs exhibit significantly better cultural knowledge recall in such challenging tasks compared to encoder-only and encoder-decoder LLMs, highlighting their superior capacity as knowledge bases Petroni et al. (2019). Besides, results across all models (Avg. Column) indicate that the groups from the U.S. and India, whose official languages include English, perform the best generally. They consistently rank in the top 3 in both automatic evaluation metrics. This pattern is especially pronounced in the results for encoder-only LLMs and encoder-decoder LLMs. Apart from the Iran group [^7] See detailed analysis in Appendix E., these two cultural groups also demonstrate the best performance in decoder-only LLMs. Especially, for the ‘dumb’ baseline, the U.S. probing performance is not the best, suggesting that the data distribution does not influence the LLM’s cultural judgment capabilities. This further strengthens the persuasiveness of our experiment.
Additionally, (1) to assess the variation in model performance across different cultural groups, we calculate the Coefficient of Variation based on the absolute-match mAP metric. This serves as an indicator of the degree of cultural bias in each model’s performance. Our results show that decoder-only LLMs exhibit less cultural bias compared to other model types. Among these, the newer version, Llama3, demonstrates greater consistency in performance than Llama2. Similarly, the non-English-dominant Qwen2 LLM also shows smaller variations in cultural bias. (2) To explore the impact of different embedding methods in mWS, we use BERT Base with mean pooling to evaluate probing performance under the mWS metric. Results (see Table 8) show that the U.S. cultural group achieves the highest performance, while decoder-only LLMs exhibit significant improvement, consistent with findings using FastText. We calculate the correlation between mWS evaluations using BERT Base and FastText for each LLM. The results reveal a strong correlation for decoder-only LLMs, as shown in Table 1. (3) To explore the correlation between training data size and cultural probing performance, we use country dish counts as a proxy for Wikidata’s cultural data size and calculate their correlation with probing performance. The correlation results are show as Table 2. Based on the probing performance evaluated by the two automated metrics, LLMs like Bb, B1, and T5 consistently exhibit little correlations with cultural data size. In contrast, decoder-only LLMs show almost no correlations, indicating limited sensitivity to cultural data size as represented by dish counts. Therefore, in more advanced LLMs, cultural competency isn’t directly tied to the training sample size of specific cultural groups.
Figure 2: Comparative probing results on FmLAMA-en: incorporating cultural information about the origin of dishes into English prompts can enhance the probing of cultural knowledge. Significance levels are indicated by * for $p<0.1$, **∗ ∗ for $p<0.05$, and ***∗ ∗ ∗ for $p<0.01$.
5.3 Manual evaluation results
Table 1 shows that Llama3 achieve the best probing performance, so we conduct manual evaluation on its probing results. Table 3 displays the results of the GPT-4o simulator and human evaluations, showing a high Pearson correlation of 0.88. Both evaluation methods consistently show that the highest MES scores come from English-speaking cultural groups, including the U.S., U.K., and India, with the U.S. performing especially well in both manual evaluations. This aligns closely with the automatic evaluation results, reinforcing the argument that LLMs tend to exhibit cultural bias toward these groups.
Figure 3: Comparison of probing results on Llama3 with prompts in different languages, showing that the English prompt exhibits the best performance.
Figure 4: Probing results using code-switching prompts on Qwen2 and LLaMA3.
Upon reviewing the evaluation results for the GPT-4o simulator, we find that it treats synonymous expressions as Direct Matches, e.g., cooked rice vs. gohan (the Japanese term) and cauliflower vs. Brassica oleracea var. botrytis (the botanical term). Additionally, reference labels from Wikidata often exhibit inconsistencies, with terms like "shrimp or prawn" and "fish as food" marked as ingredients. Furthermore, GPT-4o flags pre-made ingredients as Substitutable; for example, flour in Garganelli with pre-made pasta. For human evaluators, evaluation results may vary due to individual subjective biases. The average Inter-Rater Reliability (IRR) across country groups, measured by Fleiss’ kappa, is 0.57, [^8] IRR and the average evaluation time for each country group are shown in Appendix C. indicating moderate agreement. While this indicates acceptable reliability for subjective tasks, it also highlights that evaluators from the same cultural background may still show variation in their judgments. Although GPT-4o can partially simulate human evaluations, its hallucination issues Ji et al. (2023) raise concerns about the reliability of its assessments. These observations underscore the inherent challenges of using automatic metrics for culturally-related tasks.
6 Analysis: Designing Better Prompts
We examine the impact of different prompt settings on cross-cultural knowledge exploration, focusing on cultural background references, multilingual prompts, and code-switching scenarios.
Cultural background analysis.
Using the English prompts from Figure 9 as a basis, we incorporate information about the dish country of origin into the probing prompts for each specific dish (as described in §4.2). Figure 2 presents the comparative probing results on FmLAMA-en across different LLMs, considering the inclusion of cultural information mentions in the prompts. We find that English prompts with cultural information achieve higher probing scores in capturing LLMs’ knowledge within the food domain. This suggests the importance of emphasizing the cultural background when utilizing LLMs, especially in the exploration of culture-related topics.
Multilingual prompt analysis.
We utilize prompts in six different languages to conduct knowledge probing on the multilingual LLM, Llama3.999The probing results comparison on other multilingual LLMs are shown in Appendix D.2. To ensure a fair comparison of probing results across different language prompts, we filter a sub-dataset where all food instances have both subject and ingredient labels available in all the languages involved. Figure 3 displays probing results on the filtered sub-dataset with prompts in different languages. [^10] Some Chinese and Korean objects are missing Fasttext vectors. Therefore, mWS cannot be calculated for them. The English prompt performs significantly better than prompts in other languages across all cultural groups, probably due to more training data.
Code-switching analysis.
Figure 4 presents the probing results using code-switching prompts, for each pair out of six languages, applied to the two best-performing decoder-only LLMs, Qwen2 and Llama3. When English is the main language and the dish subject is presented in another language, probing performance significantly declines but still outperforms probing conducted solely in other languages. This suggests that the instruction language is important, not just the language of the dish name. Furthermore, when the main language of the prompt is non-English and the subject language is English, performance does not significantly decline and even shows a slight improvement. This further emphasizes the dominant role of English in multilingual LLMs. The detailed probing results of each cultural group and other multilingual LLMs are in shown Appendix D.3.
Country Top 5 Most Falsely Predictions (Llama3)
Italy pasta, bread, meat, egg, pork
Japan rice, fish, soy, tofu, chicken
France butter, cream, meat, caramel, bread
Russia meat, beef, cabbage, pork, beet
(a) Coarse-grained confusions
Model Top 5 Most Frequent Predictions
BERT dried meat, ham, beef, pork, fish
T5 vinaigrette, scrambled eggs, buckwheat flour, ricotta, sultana
Llama2 rice, meat, bread, white meat, egg
Llama3 rice, meat, sugar, chicken, bread
(b) Fine-grained confusions
Table 4: Error Analysis: Coarse and fine-grained confusions in ingredient predictions by various cultural groups and models.
7 Error Analysis
A manual inspection of a data sample [^11] Conducted by one of the authors on 500 predictions by Llama3, Llama2, Bert-base-uncased and T5-base. followed by an automated evaluation on the complete data reveals recurring mistake patterns: (1) Coarse-grained confusions (foreign ingredients, e.g., coffee in a Chinese dish), where the main issue lies in the model’s lack of cultural attribute awareness for both the dish and its ingredients. (2) Fine-grained confusions (wrong local ingredients, e.g., Chinese wine in Mapo tofu), and (3) Inconsistent confusions, where the model predicts different incorrect ingredients across prompts for the same dish.
Table 4 illustrates the described error patterns. In the results for Llama3 in English, $47\%$ of the incorrect predictions for Italian dishes are the ingredient pasta, $45\%$ of the incorrect predictions for French dishes are butter, and $40\%$ of the incorrect predictions for Japanese dishes are tofu. These are examples of fine-grained confusions. Additionally, rice, wheat and flour account for $11\%$ of the Llama3’s total predictions, regardless of the dish’s origin. This is a coarse confusion, where the model often predicts the same foreign ingredient.
These errors reveal a lack of cultural knowledge about the dishes and suggest different types of guessing. While coarse and inconsistent confusions seem to reflect more general guesses, possibly influenced by the frequency of certain ingredients in the training data, fine-grained confusions indicate that the model has some knowledge of the dish’s origin and is attempting to guess a local ingredient accordingly (i.e. overfitting to the context). Specifically, for fine-grained errors, even when the cultural attributes of candidate ingredients align with the dish, LLMs still make incorrect predictions. While many studies have shown that LLMs can function as knowledge bases Nguyen et al. (2024); Pan et al. (2024) and memorize factual information Fierro et al. (2024a, b), the familiarity with specific entities is crucial in shaping their predictive expectations Du et al. (2024).
8 Conclusion
This study presents an automated method for generating extensive cultural knowledge datasets, exemplified by the creation of FmLAMA, a diverse, food-centric dataset that spans multiple cultures and languages. We introduce novel metrics for cultural knowledge evaluation in LLMs, emphasizing the influence of cultural context and language in the probing process. Our findings reveal a predominant bias towards American culture in LLMs when using English prompts, a bias that diminishes with prompts in other languages. Interestingly, incorporating explicit cultural cues in prompts enhances LLMs’ cultural knowledge access. The study also highlights the scarcity of culturally diverse knowledge across languages, pointing to a potential root of observed biases in LLMs.
9 Limitations
While this study provides valuable insights into cross-cultural knowledge probing in LLMs, it is essential to acknowledge several limitations. Firstly, the food domain knowledge dataset utilized in this research is sourced from Wikidata, which may not offer comprehensive coverage. For example, the dish soy sauce chicken may only include the ingredient chicken meat while lacking the inclusion of soy sauce. Moreover, ingredient descriptions are not always detailed. For instance, the Wikidata gold label might be oil when the recipe requires a specific type of oil, such as sesame oil. This inconsistency underscores the motivation behind our mWS metric. Furthermore, aside from well-known dishes, certain recipes lack standardization and may vary depending on individual preferences and cooking styles, posing challenges to precise probing. Additionally, the fuzzy-match metric mWS, introduced in this study, relies on Fasttext for obtaining object representation vectors. However, for certain objects in Chinese and Korean, zero vectors may result, rendering similarity calculation impossible. Furthermore, we employ manually crafted templates in this paper. However, research has shown that sampling templates from large corpora can also enhance knowledge-probing evaluation. This aspect is deferred to future work. Despite our endeavors to construct comprehensive multilingual and multicultural knowledge repositories, the availability of aligned cross-cultural knowledge remains limited in multilingual settings. This constraint presents challenges in exploring the interaction between language and culture.
ACKNOWLEDGMENTS
This research is supported by the project of Shenzhen Science and Technology Research Fund (Fundamental Research Key Project Grant No. JCYJ20220818103001002), Shenzhen Science and Technology Program (Grant No. ZDSYS20230626091302006), Key Project of Shenzhen Higher Education Stability Support Program (Grant No. 2024SC0009).
Appendix A FmLAMA details
Datasets Format Topic Construction method Size
GeoMLama  Yin et al. (2022) Manual Template Geo-Diverse Concept Manually curated 3125
FORK Palta and Rudinger (2023) CommonsenseQA Culinary culture Manually curated 184
StereoKG  Deshpande et al. (2022) Triplet knowledge Stereotypes about religion and ethnicity Automatically constructed 4722
Candle  Nguyen et al. (2023) Sentences Several cultural facets (food, drinks, clothing, traditions, rituals, behaviors) Automatically constructed 47360
FmLAMA (ours) Triplet knowledge Food domain Automatically constructed 33600
Table 5: Comparison of cultural knowledge datasets. Size is in number of instances.
A.1 Examples
Figure 5 illustrates examples from dataset FmLAMA. Each dish instance in FmLAMA is defined as: $\left(url,na,cou,la,pa,\left[ma,im\right]\right)$, the elements in []delimited-[][$\cdot$][ ⋅ ] indicate optional:
  • $url$: the link in Wikidata;
  • $na$: the name of the dish;
  • $cou$: the country of origin of the dish;
  • $la$: the language used in this entry;
  • $pa$: the ingredients of this dish;
  • $ma$: the material used in the dish;
  • $im$: the image of this dish.
Particularly, in each instance, $la=\mathrm{LANG}\left(na\right)=\mathrm{LANG}\left(pa\right)=\mathrm{LANG}\left(ma\right)$, which indicates the descriptive language for all attributes remains consistent. For a dish with the same $url$, there may be several instances with different languages. Table 5 contrasts our dataset with prior cultural knowledge collections.
A.2 Statistics
There are a total of 33,600 dishes in FmLAMA. We count the number of dishes corresponding to each language and country in the dataset, as shown in Figures 6 and 7. Figure 6(a) shows the top 20 languages ranked by the number of dishes, with English (en) having the highest count at 2804, followed by Spanish (es), French (fr), and Japanese (ja). Figure 6(b) displays the distribution of dish counts across different intervals for all languages, indicating that most languages have fewer than 200 dishes. Similarly, Figure 7(a) shows the top 20 countries ranked by the number of dishes, with Italy having the highest count at 2975, followed by France, America, and Japan. Figure 7(b) shows the distribution of dish counts across various intervals for all countries, indicating that only 8 countries have more than 200 dishes. Figure 8 shows the dish statistics by ingredient count. The majority of dishes contain only one ingredient, totaling 18,546. Dishes with two ingredients are the second most common, with a count of 6,426. The counts decrease as the number of ingredients increases, with dishes containing eleven ingredients being the least common, at 32.
(a) The Top20 languages ranked by the dishes count
(b) Languages distributed across various dish count ranges
Figure 6: Dish statistics by languages
(a) The Top20 countries ranked by the dishes count
(b) Countries distributed across various dish count ranges
Figure 7: Dish statistics by countries
Figure 8: Dish statistics by ingredient count
Appendix B Probing Templates
Figure 9: Probing templates in six involved languages, with [X] representing the subject and [Y] indicating the object that can be substituted.
B.1 Prompt illustration
Probing templates in six involved languages are shown in Figure 9, in which [X] represents the subject (dish) and [Y] indicates the object (ingredient).
B.2 Language comparison
B.2.1 Geographic differences
In this section, we discuss the representative countries where each selected language is spoken and their geographic distribution: (1) English: Mainly spoken in U.S., Canada, Australia, Guyana, etc., spanning North and South America, Europe, Australia, and Africa. (2) Chinese: Mainly spoken in China, Singapore, etc., covering huge area of Asia. (3) Arabic: Mainly spoken in Saudi Arabia, Egypt, United Arab Emirates, etc., covering the Arabian Peninsula, North Africa, and some sub-Saharan African countries. (4) Korean: Mainly spoken in South Korea, North Korea. (5) Russian: Mainly spoken in Russia, Belarus, Kazakhstan, etc., covering Eastern Europe, Central Asia, and Northern Asia. (6) Hebrew: Mainly spoken in Israel, with Hebrew-speaking communities worldwide. These languages represent cultural diversity, being spoken on different continents by groups with rich and distinct cultural backgrounds, demonstrating the broad geographic and cultural diversity represented by these languages.
B.2.2 Grammatical differences
In addition to geographical differences in usage, these six languages also display the following grammatical and reading distinctions: (1) Word Order: English and Chinese follow an Subject-Verb-Object word order, Arabic typically uses Verb-Subject-Object, Korean follows Subject-Object-Verb, while Russian and Hebrew are relatively flexible but generally adhere to Subject-Verb-Object. (2) Tense: English, Arabic, Russian, and Hebrew possess complex tense systems. Chinese lacks a strict tense structure, whereas Korean expresses tense through verb conjugation. (3) Gender: Arabic, Russian, and Hebrew have grammatical gender, while English distinguishes gender only in pronouns. Chinese and Korean do not have grammatical gender. (4) Articles: English, Arabic, and Hebrew have an article system, whereas Chinese, Korean, and Russian do not use articles. (5) Reading Direction: English, modern Chinese, modern Korean, and Russian are read from left to right, while Arabic and Hebrew are read from right to left. Traditional Chinese and traditional Korean are read vertically from top to bottom, with columns arranged from right to left. (6) Case System: Arabic and Russian have complex case systems, while Korean employs particles to indicate grammatical relations. English, Chinese, and Hebrew do not use case inflections.
Appendix C Manual Evaluation Setup
C.1 GPT-4o simulator
The evaluation prompts for the GPT-4o simulator are shown in Table 6.
C.2 Human evaluator
Figure 10 shows the human evaluation website. The Inter-Rater Reliability (IRR) and the average evaluation time for the three evaluators from each country group are provided in Table 7.
Instructions
You are a professional evaluator. I will provide you with a dish, its reference ingredient label sourced from Wikidata, and a list of predicted ingredients. Your task is to evaluate whether each predicted ingredient is appropriate for the respective dish.
### Evaluation Criteria:
For each predicted ingredient, mark it as Correct, Maybe, or Incorrect using the following guidelines:
Correct: The ingredient meets one of these: Direct Match: It is explicitly listed in the reference ingredient label from Wikidata. Substitutability: It can replace a specific ingredient in the reference label during cooking. Missing Traditional Ingredient: It is traditionally or commonly used in the dish, but not listed in the reference label. Maybe: The ingredient could be used in some variations of the dish, but its use is uncommon or ambiguous. Incorrect: The ingredient is rarely or never used in the dish. ### Output Format:
Please directly return the evaluation as a Python dictionary with JSON format where the key is the predicted ingredient, and the value is a tuple consisting of:
1. The evaluation ("Correct", "Maybe", "Incorrect")
2. The reason ("Direct Match", "Substitutability", "Missing Traditional Ingredient", or a brief explanation)
### Example:
Input:
Dish: Spaghetti Bolognese
Reference ingredient label: spaghetti, ground beef, onion, garlic, Carrot
Predicted ingredients: spaghetti, basil, chicken, tomato, beef
Output
    ’ground beef’: (’Correct’, ’Substitutability’),
    ’meat’: (’Maybe’, ’Ambiguous’),
    ’beef’: (’Correct’, ’Substitutability’),
    ’pasta’: (’Incorrect’, ’Pasta is not part of the sauce itself’),
    ’turkey meat’: (’Maybe’, ’It can be used in some variations’),
Table 6: GPT-4o simulator: Evaluation Instructions
Figure 10: Human evaluation website
Group Dishes IRR Time(m)
Italy 215 0.60 74.93
U.S. 285 0.48 104.25
Turkey 98 0.78 32.05
Japan 186 0.64 49.25
France 175 0.63 71.83
U.K. 83 0.46 30.33
Mexico 57 0.44 22.25
India 132 0.58 40.25
Germany 57 0.69 20.03
China 97 0.59 48.50
Iran 21 0.61 13,33
Greece 21 0.35 11.30
Spain 95 0.57 41.50
Russia 27 0.51 26.50
Mean 111 0.57 -
Table 7: Human evaluator: IRR and average evaluation time for the three evaluators from each country group.
Appendix D Supplementary Experiments and Analyses
D.1 Probing results in each FmLAMA-la
In addition to the multilingual LLMs discussed in §5, configure probing for monolingual, encoder-only LLMs in Arabic, Russian, Korean, and Chinese, including asafaya/bert-base-arabic Safaya et al. (2020), DeepPavlov/rubert-base-cased Kuratov and Arkhipov (2019), kykim/bert-kor-base, klue/bert-base Park et al. (2021) and bert-base-chinese.
The probing results with prompts in the other five languages (Arabic, Chinese, Hebrew, Korean, and Russian) on the corresponding filtered sub-datasets (FmLAMA-ar, FmLAMA-zh, FmLAMA-he, FmLAMA-ko, and FmLAMA-ru) are depicted in Table 9101112, and 13, respectively. Because certain objects in Chinese and Korean have representation vectors that result in all zeros when obtained through Fasttext, calculating cosine similarity was not feasible. Consequently, mWS evaluation was not conducted for prompts in Chinese and Korean prompts.
Origin Encoder-only LLMs Encoder-Decoder LLMs Decoder-only LLMs Avg.
Bb Bl mB mT5 T5 Qwen2 Llama2 Llama3
Italy 0.7393±0.01 0.7425±0.01 0.7438±0.00 0.6865±0.02 0.6579±0.01 0.7749±0.01 0.7658±0.01 0.7869±0.01 0.7372
U.S. 0.7748±0.01 0.7763±0.01 0.7815±0.01 0.7071±0.02 0.7178±0.03 0.7968±0.01 0.7839±0.02 0.8107±0.01 0.7686
Turkey 0.7357±0.01 0.7461±0.01 0.7545±0.01 0.7015±0.02 0.6703±0.01 0.7681±0.01 0.7571±0.01 0.7830±0.01 0.7395
Japan 0.7113±0.00 0.7144±0.00 0.7218±0.01 0.6788±0.02 0.6442±0.01 0.7446±0.00 0.7446±0.01 0.7516±0.00 0.7139
France 0.7398±0.01 0.7520±0.02 0.7513±0.01 0.6755±0.02 0.6642±0.01 0.7941±0.01 0.7796±0.01 0.7846±0.00 0.7426
U.K. 0.7460±0.01 0.7500±0.01 0.7550±0.00 0.6697±0.02 0.6745±0.02 0.7678±0.01 0.7701±0.01 0.7790±0.01 0.7390
Mexico 0.7273±0.01 0.7313±0.01 0.7291±0.01 0.7171±0.01 0.6734±0.01 0.7605±0.01 0.7752±0.02 0.7852±0.02 0.7374
India 0.7477±0.01 0.7523±0.00 0.7568±0.00 0.6998±0.01 0.6619±0.01 0.7867±0.00 0.7944±0.01 0.7935±0.01 0.7491
Germany 0.7348±0.01 0.7305±0.03 0.7493±0.00 0.6755±0.01 0.6535±0.01 0.7835±0.01 0.7769±0.02 0.7849±0.01 0.7361
China 0.7273±0.01 0.7319±0.01 0.7317±0.01 0.7024±0.01 0.7054±0.01 0.7698±0.01 0.7593±0.01 0.7793±0.01 0.7384
Iran 0.7064±0.01 0.7013±0.01 0.7310±0.01 0.7042±0.02 0.6686±0.01 0.7905±0.03 0.7914±0.04 0.8026±0.01 0.7370
Greece 0.7488±0.01 0.7618±0.02 0.7566±0.02 0.7190±0.03 0.6369±0.00 0.8137±0.04 0.7739±0.01 0.8247±0.01 0.7544
Spain 0.7402±0.00 0.7465±0.01 0.7472±0.00 0.6883±0.01 0.6391±0.01 0.7748±0.01 0.7663±0.01 0.7765±0.01 0.7349
Russia 0.7467±0.01 0.7571±0.01 0.7573±0.00 0.7216±0.02 0.6710±0.01 0.7716±0.01 0.7680±0.01 0.7721±0.00 0.7457
ALL 0.7419±0.00 0.7466±0.01 0.7507±0.00 0.6924±0.01 0.6723±0.01 0.7778±0.01 0.7711±0.01 0.7859±0.00 0.7423
Corr. 0.35 0.59 0.38 0.83 0.57 0.86 0.71 0.96 -
Table 8: Probing performance comparison with English prompts and FmLAMA-ar sub-dataset. The results are evaluated using mWS with BERT base. Corr. denotes the correlation of mWS evaluations using BERT Base and FastText for each LLM, which reveal a strong correlation on decoder-only LLMs,
Origin Count Bb-ar mB mT5 Qwen2 Llama2 Llama3 Avg.
Italy 76 (19.2%) 2.56±0.69 3.07±1.02 5.12±1.20 5.13±3.42 3.30±1.13 4.05±0.81 4.03
U.S. 54 (13.7%) 5.29±1.18 7.22±3.91 5.68±1.53 9.41±4.12 3.93±1.00 5.95±2.38 5.09
Turkey 52 (13.2%) 4.07±0.88 6.66±3.09 3.30±0.53 9.20±5.71 3.33±1.04 9.54±3.48 5.47
Japan 44 (11.11%) 2.69±1.50 2.56±0.50 2.13±0.78 9.91±3.73 3.75±1.12 5.43±1.62 5.20
France 37 (9.4%) 7.51±2.43 6.50±4.55 4.51±1.21 7.16±3.23 5.71±0.82 4.63±1.77 4.70
U.K. 26 (6.6%) 5.80±2.15 9.17±6.85 6.72±1.67 11.77±6.35 5.72±4.86 8.62±6.52 5.75
Mexico 20 (4.6%) 3.52±1.57 7.37±1.01 3.43±0.61 3.60±0.55 2.38±0.48 3.72±1.68 2.82
India 18 (3.2%) 4.39±1.09 8.86±5.19 3.08±1.17 10.82±9.44 3.04±1.56 3.03±1.38 7.29
Germany 13 (3.3%) 7.39±2.52 10.65±4.95 5.87±2.04 9.53±4.72 2.75±2.12 7.67±3.29 4.69
China 13 (3.3%) 2.07±0.53 4.86±4.36 2.43±0.93 6.79±2.17 2.75±0.70 6.54±2.51 4.58
Iran 11 (2.8%) 9.30±4.85 8.19±5.41 7.38±2.56 11.23±7.82 6.31±2.85 6.67±5.02 5.90
Greece 11 (2.8%) 3.12±0.98 5.21±3.26 4.25±1.23 9.25±4.78 3.13±1.09 4.54±3.95 4.63
Spain 10 (2.5%) 5.84±1.24 4.13±2.32 8.01±1.35 12.74±3.84 3.41±1.18 6.90±4.08 7.34
Russia 10 (2.5%) 3.62±0.80 3.13±0.79 2.48±0.26 4.20±0.98 4.88±2.78 4.76±0.94 4.34
ALL 395 (100.0%) 4.41±0.35 5.80±2.56 4.48±0.60 8.24±3.65 3.85±1.00 5.85±1.47 4.96
(a) Performance results evaluated on mAP (%).
Origin Bb-ar mB mT5 Qwen2 Llama2 Llama3 Avg.
Italy 0.3280±0.01 0.2259±0.02 0.3547±0.05 0.3213±0.06 0.2252±0.05 0.2802±0.05 0.2892
U.S. 0.3355±0.02 0.2746±0.05 0.3581±0.03 0.3261±0.05 0.2419±0.03 0.3194±0.02 0.3093
Turkey 0.3141±0.01 0.2434±0.04 0.3154±0.03 0.3111±0.06 0.2210±0.03 0.3164±0.06 0.2869
Japan 0.2876±0.02 0.2411±0.04 0.3150±0.05 0.3030±0.05 0.1834±0.04 0.2553±0.04 0.2642
France 0.3599±0.02 0.2984±0.04 0.3462±0.03 0.3129±0.04 0.2608±0.02 0.2844±0.01 0.3104
U.K. 0.3450±0.02 0.3118±0.08 0.3651±0.04 0.3450±0.06 0.2406±0.07 0.3435±0.05 0.3252
Mexico 0.3098±0.02 0.2891±0.04 0.3335±0.07 0.2687±0.03 0.2163±0.02 0.2665±0.02 0.2807
India 0.2850±0.03 0.2474±0.07 0.3051±0.03 0.3355±0.09 0.2217±0.05 0.2455±0.03 0.2734
Germany 0.3300±0.02 0.2859±0.07 0.3530±0.04 0.2772±0.06 0.2046±0.04 0.3172±0.05 0.2946
China 0.2938±0.02 0.2836±0.06 0.3255±0.07 0.2889±0.03 0.1907±0.04 0.2774±0.04 0.2767
Iran 0.3453±0.04 0.2644±0.06 0.3634±0.03 0.3488±0.04 0.2405±0.06 0.3257±0.02 0.3147
Greece 0.3241±0.02 0.2559±0.04 0.3502±0.05 0.3358±0.06 0.2325±0.03 0.3052±0.04 0.3006
Spain 0.3659±0.02 0.3213±0.03 0.4071±0.06 0.3626±0.06 0.2869±0.05 0.3576±0.05 0.3502
Russia 0.3323±0.02 0.2772±0.02 0.3347±0.02 0.3451±0.02 0.2964±0.03 0.3145±0.00 0.3167
ALL 0.3243±0.01 0.2625±0.03 0.3421±0.04 0.3176±0.05 0.2282±0.04 0.2957±0.03 0.2951
(b) Performance results evaluated on mWS.
Table 9: Probing performance comparison with Arabic prompts and FmLAMA-ar sub-dataset.
Origin Count Bb-zh mB mT5 Qwen2 Llama2 Llama3 Avg.
Italy 49 (8.7%) 17.18±1.67 13.07±4.92 10.55±0.84 24.54±2.87 9.32±4.48 13.58±4.08 14.71
U.S. 101 (17.9%) 19.32±1.47 13.79±3.56 12.37±1.09 26.50±2.94 12.91±3.61 13.17±6.45 16.34
Turkey 12 (2.1%) 8.92±3.86 6.30±2.12 5.22±0.18 18.16±3.49 8.18±3.58 11.00±6.39 9.63
Japan 114 (20.2%) 15.79±2.32 11.18±4.41 7.85±0.60 17.67±2.73 11.15±2.66 11.78±4.54 12.57
France 70 (12.4%) 19.45±4.13 16.04±5.31 9.14±0.98 20.96±1.45 13.36±3.62 14.71±5.89 15.61
U.K. 38 (6.7%) 27.88±2.65 19.55±7.36 10.64±0.68 28.43±4.16 18.02±8.33 21.46±6.75 21.00
Mexico 19 (3.4%) 21.12±0.86 13.40±3.51 12.05±0.79 16.25±4.70 10.60±1.32 8.21±3.83 13.61
India 24 (4.3%) 5.47±1.65 5.02±2.68 1.13±0.17 12.37±3.79 4.57±2.20 6.58±5.13 5.86
Germany 16 (2.8%) 10.31±2.53 12.46±3.54 6.81±1.75 10.41±2.12 4.08±2.38 7.62±4.63 8.62
China 87 (15.4%) 22.55±4.86 18.32±6.75 7.88±0.91 33.48±6.15 13.83±3.72 19.96±9.42 19.34
Iran 7 (1.2%) 7.72±1.44 8.19±4.58 3.54±0.20 8.39±2.08 3.40±0.38 8.82±3.50 6.68
Greece 7 (1.2%) 4.11±0.68 2.88±1.08 5.43±0.43 14.81±3.05 1.40±0.37 3.46±1.23 5.35
Spain 12 (2.1%) 13.00±1.83 10.21±1.58 12.72±1.17 21.02±0.93 5.53±3.77 12.41±6.05 12.48
Russia 8 (1.4%) 5.58±0.93 3.98±0.76 2.29±0.38 5.89±1.09 4.06±2.84 5.14±1.07 4.49
ALL 564 (100%) 17.84±2.37 13.56±4.36 8.96±0.44 22.71±2.37 11.46±3.32 13.76±5.56 14.71
Table 10: Probing performance comparison with Chinese prompts and FmLAMA-zh sub-dataset.
Origin Count mB mT5 Qwen2 Llama2 Llama3 Average
Italy 59 (18.6%) 4.18±2.54 4.30±1.49 7.78±2.37 3.97±1.47 6.29±3.14 5.30
U.S. 67 (21.1%) 8.62±7.22 8.13±2.02 9.09±4.52 3.62±0.81 10.17±4.44 7.93
Turkey 12 (3.9%) 8.99±8.65 6.33±4.01 9.21±5.92 2.93±1.18 10.68±6.33 7.63
Japan 19 (6.0%) 6.69±5.23 5.99±3.64 6.60±4.19 6.92±3.85 4.39±3.12 6.12
France 65 (20.5%) 5.43±3.01 4.54±0.92 6.40±2.74 2.72±0.54 5.50±1.47 4.92
U.K. 23 (7.3%) 9.43±6.79 6.95±2.26 9.79±5.06 3.97±2.49 10.08±5.66 8.04
Mexico 11 (3.5%) 2.22±0.75 2.26±0.74 4.33±2.44 2.87±1.99 2.14±2.01 2.76
India 18 (5.7%) 8.12±3.72 3.31±1.07 2.80±2.77 3.26±1.95 4.73±2.46 4.44
Germany 11 (3.5%) 7.57±5.98 14.84±6.23 8.59±3.46 5.92±1.83 9.38±5.42 9.26
China 9 (2.8%) 3.61±3.34 2.12±0.78 3.75±3.62 1.64±1.29 4.05±1.62 3.03
Iran 6 (1.9%) 8.58±5.37 5.05±2.90 19.40±9.90 2.95±0.57 20.63±10.96 11.32
Greece 6 (1.9%) 2.76±0.60 4.31±0.49 12.96±9.55 7.67±2.85 3.34±2.19 6.21
Spain 7 (2.2%) 4.58±1.81 7.02±1.87 3.86±1.66 4.04±1.62 2.82±1.26 4.46
Russia 4 (1.3%) 4.31±3.65 3.35±0.07 2.10±0.28 2.56±0.46 2.56±0.44 2.98
ALL 317 (100%) 6.41±4.30 5.77±1.59 7.57±2.66 3.74±0.85 7.17±2.63 6.13
(a) Performance results evaluated on mAP (%).
Origin mB mT5 Qwen2 Llama2 Llama3 Average
Italy 0.3102±0.03 0.3320±0.03 0.3197±0.03 0.2973±0.04 0.3265±0.05 0.3171
U.S. 0.3435±0.06 0.3626±0.03 0.3524±0.03 0.3042±0.03 0.3517±0.04 0.3429
Turkey 0.2976±0.06 0.3225±0.05 0.3110±0.04 0.2621±0.05 0.3296±0.10 0.3046
Japan 0.3134±0.05 0.3167±0.04 0.3114±0.05 0.2839±0.03 0.2899±0.02 0.3031
France 0.3082±0.03 0.3509±0.04 0.3334±0.02 0.2838±0.04 0.3134±0.03 0.3179
U.K. 0.3400±0.06 0.3418±0.04 0.3414±0.04 0.2930±0.02 0.3167±0.06 0.3266
Mexico 0.2959±0.05 {0.3829±0.06} 0.3420±0.06 0.2847±0.06 0.2865±0.04 0.3184
India {0.3614±0.05} 0.3437±0.06 0.3354±0.01 0.3133±0.04 0.3192±0.05 0.3346
Germany {0.3595±0.04} 0.3763±0.02 {0.3756±0.03} {0.3339±0.05} {0.3586±0.06} 0.3608
China 0.2863±0.03 0.3131±0.06 0.2714±0.06 0.2683±0.03 0.2432±0.04 0.2765
Iran 0.3435±0.05 0.3521±0.07 {0.4569±0.07} 0.2902±0.02 {0.4416±0.09} 0.3769
Greece 0.3014±0.02 0.3423±0.02 0.3722±0.08 0.3332±0.03 0.2951±0.04 0.3288
Spain 0.3493±0.03 {0.4229±0.07} 0.3547±0.02 {0.3635±0.06} 0.3260±0.03 0.3633
Russia 0.3276±0.03 0.3748±0.06 0.3319±0.04 0.3308±0.01 0.3179±0.04 0.3366
ALL 0.3237±0.04 0.3484±0.03 0.3369±0.02 0.2969±0.03 0.3248±0.04 0.3261
(b) Performance results evaluated on mWS.
Table 11: Probing performance comparison with Hebrew prompts and FmLAMA-he sub-dataset.
Origin Count Bb-ky Bb-kl mB-u mT5 Qwen2 Llama2 Llama3 Avg.
Italy 88 (16.7%) 6.90±1.57 7.75±2.34 3.85±2.54 6.02±0.64 3.50±1.10 1.19±0.22 2.23±0.75 4.49
U.S. 64 (12.2%) 11.37±2.66 10.87±3.86 6.32±5.20 10.15±0.34 2.46±0.69 1.68±0.12 2.95±0.93 6.54
Turkey 19 (3.6%) 2.75±2.47 4.13±2.62 4.07±3.17 1.82±0.44 3.86±1.00 2.16±0.12 7.47±1.45 3.75
Japan 101 (19.2%) 7.34±2.12 7.74±1.86 1.93±1.24 5.25±0.99 6.56±5.34 2.19±0.22 2.62±1.15 4.80
France 67 (12.7%) 2.33±0.43 3.09±0.44 3.15±0.17 2.79±0.30 3.10±0.68 1.40±0.09 2.41±1.27 2.61
U.K. 27 (5.1%) 14.07±2.83 15.89±4.43 6.00±4.42 7.55±1.34 3.21±1.02 1.79±0.31 6.89±5.70 7.91
Mexico 13 (2.5%) 5.18±3.69 8.37±2.85 4.89±3.29 2.31±0.40 7.12±3.13 2.75±1.99 4.18±1.61 4.97
India 32 (6.1%) 7.98±2.29 5.74±1.32 3.28±2.09 2.96±0.61 7.69±6.42 2.01±0.32 2.36±0.67 4.57
Germany 15 (2.9%) 2.50±0.94 5.08±1.25 2.42±1.68 3.20±1.18 2.63±0.95 1.33±0.48 1.91±0.21 2.72
China 54 (10.3%) 9.02±2.76 10.12±3.04 2.36±2.22 4.05±1.50 4.43±1.74 2.16±0.48 3.06±1.48 5.03
Iran 8 (1.5%) 7.99±5.23 3.76±1.41 1.62±0.67 1.52±0.24 1.50±0.14 2.87±2.49 2.12±1.83 3.05
Greece 9 (1.7%) 2.32±0.73 6.76±3.73 2.74±0.76 5.76±4.22 4.59±1.31 3.67±2.94 3.35±1.87 4.17
Spain 21 (4.0%) 4.05±2.25 5.90±2.70 1.79±1.24 5.82±1.33 3.99±1.38 2.50±0.97 3.85±0.99 3.99
Russia 8 (1.5%) 3.70±0.99 4.22±1.83 2.68±1.12 2.12±0.05 3.01±1.41 1.72±0.20 3.07±1.46 2.93
ALL 526 (100%) 7.05±1.47 7.68±2.00 3.45±2.10 5.19±0.42 4.32±1.76 1.86±0.18 3.07±0.95 4.66
Table 12: Probing performance comparison on mAP (%) with Korean prompts and FmLAMA-ko sub-dataset.
Origin Count Bb-ru mB mT5 Qwen2 Llama2 Llama3 Avg.
Italy 101 (15.9%) 4.29±2.04 6.57±3.51 4.52±1.27 7.99±2.61 4.30±2.32 10.61±2.60 6.38
U.S. 79 (12.5%) 9.76±4.95 9.99±5.25 5.51±0.71 12.01±3.53 9.04±4.33 11.97±3.75 9.71
Turkey 28 (4.4%) 5.75±2.12 9.46±1.68 3.58±1.61 11.08±8.01 10.49±2.90 16.37±2.97 9.46
Japan 68 (10.7%) 7.89±4.76 7.13±2.34 1.76±0.56 6.10±2.74 8.50±5.79 14.11±0.74 7.58
France 93 (14.7%) 7.08±3.04 6.83±2.85 5.03±1.86 7.03±2.48 8.85±4.96 10.55±3.43 7.56
U.K. 33 (5.2%) 8.42±5.82 10.03±5.01 5.54±3.74 11.89±6.16 13.17±8.89 16.82±8.31 10.98
Mexico 22 (3.5%) 4.08±0.89 2.01±0.82 2.08±0.95 4.70±1.40 5.64±3.25 4.21±1.34 3.79
India 21 (3.3%) 11.46±9.78 19.02±2.71 4.27±3.11 16.80±6.73 12.95±5.70 22.93±2.60 14.57
Germany 45 (7.1%) 7.53±5.26 7.95±3.59 4.39±2.03 9.41±2.54 11.74±4.92 11.57±4.32 8.77
China 30 (4.7%) 9.67±4.64 11.86±3.63 3.28±2.19 10.44±4.98 12.63±6.07 12.82±4.43 10.12
Iran 13 (2.1%) 10.86±6.84 11.06±5.99 15.51±4.53 22.88±6.61 10.77±8.30 23.88±6.87 15.83
Greece 18 (2.8%) 3.45±2.42 1.74±0.43 3.38±1.24 8.56±2.57 4.91±2.85 4.77±1.28 4.47
Spain 38 (6.0%) 3.35±1.51 6.51±2.61 3.37±1.37 5.27±2.52 3.88±2.07 5.61±1.84 4.67
Russia 45 (7.1%) 4.79±2.05 3.75±0.77 2.80±0.92 5.21±2.94 5.31±2.18 5.47±2.27 4.55
ALL 634 (100%) 6.85±3.14 7.76±1.83 4.28±1.11 8.84±2.67 8.19±4.04 11.52±2.82 7.91
(a) Performance results evaluated on mAP (%).
Origin Bb-ru mB mT5 Qwen2 Llama2 Llama3 Avg.
Italy 0.3319±0.04 0.3217±0.06 0.3633±0.03 0.4073±0.04 0.3774±0.02 0.3779±0.05 0.3633
U.S. 0.3796±0.04 0.3556±0.05 0.3692±0.02 0.4393±0.03 0.3938±0.05 0.4068±0.05 0.3907
Turkey 0.3199±0.06 0.3317±0.04 0.3300±0.04 0.3966±0.06 0.4012±0.02 0.4178±0.04 0.3662
Japan 0.2955±0.03 0.2796±0.04 0.2740±0.02 0.3085±0.04 0.3110±0.05 0.3319±0.01 0.3001
France 0.3438±0.05 0.3132±0.03 0.3652±0.04 0.3875±0.02 0.3954±0.04 0.3752±0.04 0.3634
U.K. 0.3647±0.08 0.3442±0.05 0.3771±0.04 0.4327±0.04 0.4149±0.08 0.4179±0.09 0.3919
Mexico 0.3150±0.01 0.3081±0.04 0.3604±0.04 0.3688±0.03 0.3909±0.02 0.3225±0.04 0.3443
India 0.3475±0.10 0.3808±0.05 0.3062±0.03 0.4375±0.06 0.3665±0.07 0.4236±0.04 0.3770
Germany 0.3500±0.08 0.3330±0.05 0.3402±0.04 0.3968±0.03 0.4053±0.04 0.3743±0.07 0.3666
China 0.3318±0.04 0.3348±0.05 0.3276±0.03 0.3741±0.07 0.3819±0.06 0.3520±0.05 0.3504
Iran 0.3252±0.06 0.3358±0.03 0.4221±0.06 0.4916±0.06 0.4418±0.04 0.4818±0.09 0.4164
Greece 0.3360±0.05 0.3015±0.04 0.3573±0.03 0.4045±0.01 0.4012±0.03 0.3365±0.05 0.3562
Spain 0.3240±0.03 0.3283±0.04 0.3625±0.04 0.3880±0.04 0.3638±0.03 0.3313±0.04 0.3497
Russia 0.3583±0.03 0.3199±0.04 0.3899±0.03 0.3732±0.04 0.3907±0.01 0.3591±0.03 0.3652
ALL 0.3395±0.04 0.3247±0.04 0.3515±0.03 0.3940±0.03 0.3824±0.03 0.3750±0.04 0.3612
(b) Performance results evaluated on mWS.
Table 13: Probing performance comparison with Russian prompts and FmLAMA-ru sub-dataset.
D.2 Language analysis on other multilingual LLMs
Figures 111213, and 14 present a comparison of probing results for the multilingual LLMs—mBERT, mT5, Qwen2, and Llama2—using prompts in various languages on the corresponding filtered sub-datasets. Decoder-only LLMs, such as Qwen2, Llama2, and Llama3, while representing the cutting edge of current LLM technology, exhibit greater performance variation under multilingual prompt settings.
Figure 11: Comparison of probing results on mBERT with prompts in different languages,
Figure 12: Comparison of probing results on mT5 with prompts in different languages,
Figure 13: Comparison of probing results on Qwen2 with prompts in different languages,
Figure 14: Comparison of probing results on Llama2 with prompts in different languages,
D.3 Code-switching probing results in each cultural group
Figure 1528 presents a detailed comparison of the probing results for each cultural group and other multilingual LLMs.
Figure 15: Italy: Probing results of code-switching prompts.
Figure 16: U.S.: Probing results of code-switching prompts.
Figure 17: Turkey: Probing results of code-switching prompts.
Figure 18: Japan: Probing results of code-switching prompts.
Figure 19: France: Probing results of code-switching prompts.
Figure 20: U.K.: Probing results of code-switching prompts.
Figure 21: Mexico: Probing results of code-switching prompts.
Figure 22: India: Probing results of code-switching prompts.
Figure 23: Germany: Probing results of code-switching prompts.
Figure 24: China: Probing results of code-switching prompts.
Figure 25: Greece: Probing results of code-switching prompts.
Figure 26: Spain: Probing results of code-switching prompts.
Figure 27: Russia: Probing results of code-switching prompts.
Figure 28: ALL: Probing results of code-switching prompts.
Appendix E Cultural bias vs Dish popularity
Table 1 shows that some LLMs perform better in probing tasks for low-resource language groups. Specifically, the decoder-only LLM performs well in the Iran group (both mAP and mWS metric), while Greece shows decent results with both T5 and decoder-only LLMs (only mWS metric).
To understand the strong performance of the Iran group across different metrics, we analyzed the results for the 21 dishes. Table 14 shows 17 involved dishes where the predicted ingredients ranked in the top 10 (Llama3). Since the top-ranked predicted ingredients in the table are not all commonly used, we can rule out the possibility that the strong performance of the Iran group is due to the LLM’s common ingredient prediction errors. We hypothesize two possible reasons for this performance: (1) LLM Cultural Ability: Decoder-only LLMs may be more familiar with the Iran cultural group. (2) Data Bias: The dishes in the Iran group are likely well-known and representative of Iranian cuisine, so the LLM might show stronger memory performance for familiar entities Du et al. (2024).
ID Subject (dish) Involved Object (ingredients) Rank
0 pomegranate soup pomegranate 0
1 Albaloo polo rice 1
2 Kofta minced meat 0
3 Sabzi polo herb 5
4 falooda milk, vermicelli 1, 0
5 ghormeh sabzi parsley, herb 0, 4
6 Bastani milk, sugar, egg 8, 7, 6
7 zeytoon parvardeh olive 8
8 rock candy sugar 0
9 Faloodeh vermicelli 0
10 muhallebi milk, rice flour 2, 1
11 rogan josh lamb meat 1
12 ash-e doogh yogurt 0
13 Mahyawa fish 4
14 sholezard rice 1
15 Eggplant Caviar eggplant 0
16 nân-e panjere flour 6
Table 14: Top 17 dishes from the Iran group with ingredients ranked in the top 10 by Llama3, excluding common prediction errors.